### v1

# Error-Minimizing Measurements in Postselected One-Shot Symmetric Quantum State Discrimination and Acceptance as a Performance Metric

Abstract
In hypothesis testing with quantum states, given a black box containing one of the two possible states, measurement is performed to detect in favor of one of the hypotheses. In postselected hypothesis testing, a third outcome is added, corresponding to not selecting any of the hypotheses. In postselected scenario, minimum error one-shot symmetric hypothesis testing is characterized in literature conditioned on the fact that one of the selected outcomes occur. We proceed further in this direction to give the set of all possible measurements that lead to the minimum error. We have given an arbitrary error-minimizing measurement in a parametric form. Note that not selecting any of the hypotheses decimates the quality of testing. We further give an example to show that these measurements vary in quality. There is a need to discuss the quality of postselected hypothesis testing. We then characterize the quality of postselected hypothesis testing by defining a new metric acceptance and give expression of acceptance for an arbitrary error-minimizing measurement in terms of some parameters of the measurement. On the set of measurements that achieve minimum error, we have maximized the acceptance, and given an example which achieves that, thus giving an example of the best possible measurement in terms of acceptance.
Index Terms- quantum state discrimination, hypothesis testing, measurement operator, type-1 and type-2 error.
I Introduction
Quantum hypothesis testing deals with discriminating between the hypotheses corresponding to quantum properties of the nature with plethora of applications in quantum information science [2, 3, 4]. The quantum state discrimination is an example of quantum hypothesis testing which aims to make a decision based on a measurement of a given quantum object, which can be in one of the two possible states corresponding to the null and alternative hypotheses. If the decision has to be made on the basis of measurement of a single copy of the quantum object, it is called one-shot hypothesis testing and if infinitely many copies are available, it called asymptotic hypothesis testing. When the test falsely concludes the alternative hypothesis, it is called type-1 error; when the test falsely decides in the favor of the null hypothesis, it is called type-2 error. Ideally the goal is to select a measurement and a decision rule to have both type-1 and type-2 errors arbitrarily small. However, simultaneous minimization of both error is difficult. Hence, a trade-off is usually selected. For example, in asymmetric hypothesis testing, the goal is to minimize type-2 error under some constraint on the type-1 error. Another example is the symmetric hypothesis testing where the objective is to minimize the average probability of error.
In the one-shot discrimination of quantum states, the minimum possible error can be easily obtained using semi-definite programming for both symmetric and asymmetric hypothesis testing problems. In symmetric hypothesis testing, the error is characterized in a closed-form expression by the Helstrom-Holevo theorem [5, 6]. Error exponent in quantum symmetric hypothesis testing was characterized in [7, 8]. Quantum relative entropy[9] was shown to be the error exponent of asymmetric hypothesis testing in [10, 11]. Optimal measurements to achieve this error are given in [12]. In conventional quantum hypothesis testing, as described above, one of the two hypotheses is selected that concludes the presence of one of the hypotheses. Here, the key problem remains that, non-orthogonal states can not be perfectly distinguished. To avoid this problem approaches utilizing inconclusive decision rules were proposed in [13, 14, 15] which were later generalized in [16, 17, 18, 19, 20, 21, 22] focusing on maximizing confidence in the outcome of the measurement. A comprehensive survey of various quantum state discrimination strategies is given in [23, 24].
While it may not be possible to perfectly distinguish non-orthogonal states, it is possible to design measurements that include an extra outcome addition to the two outcomes corresponding to each state [25], which corresponds to not selecting any of the hypotheses (see Fig. 1). Here, the decision is only made when third outcome does not occur, hence it is termed postselected hypothesis testing. Postselected probabilities are defined as probability of any event conditioned that one of the selected outcomes is observed. The postselected error probabilities have been characterized in [25] for symmetric and asymmetric hypothesis testing, both in one-shot or asymptotic cases and derived the minimum possible error. It is also generalized to hypothesis testing for quantum channels and classical probability distributions and an example is given to show that minimum error bound is achievable in all the mentioned cases [25]. In our preliminary work [1], we define a metric acceptance signifying the probability that one of two outcomes occur and a decision is made to select one of the two hypothesis.
(a) Conventional quantum state discrimination: The measurement gives two outcomes, each corresponding to one of the possible states of the object in unknown state.
(b) Postselected quantum state discrimination: The measurement gives an addition outcome (given in red), corresponding to not selecting either of the states.
Figure 1: The unknown-state quantum object is in the state $\nu\in\{\rho,\sigma\}$. A measurement is performed to detect the state.
Motivation and Contributions: It is reasonable to wonder if finding an error minimization-measurement is sufficient in postselected hypothesis testing. Let us consider a case where there are two measurements having the same error but different probability of rejection. Essentially, even though error is same, the one having lower rejection is better. Further, while it may be possible that a measurement that minimizes the error has a high probability of having third outcome (i.e. rejection). As such measurements may practically be unusable, this brings a question if and how we can find the measurements that minimize the probability of rejection as well as the postselected error. The key contributions of present work is as follows:
  1. 1.
    We begin with deriving the condition on the measurements for it to be an error-minimizing measurement. We obtain three different conditions in three different cases depending on the prior probability of unknown state being $\rho$ (and $\sigma$). Based on the conditions, we observe that the error-minimizing measurement never makes decision in favor $\rho$ (or $\sigma$), if prior probability of unknown state being in the state $\rho$ i.e. $p_{\rho}$ is smaller (or greater) than a threshold. If it is equal to the the threshold, the error-minimizing measurement may make a decision in favor of either/both of the states. A classical example (see Example 1) is given to illustrate finding measurements from this and further showing that acceptance varies over the set of all error minimizing measurement.
  2. 2.
    We have given the set of all error-minimizing measurements in parameterized form for all the three cases. This a generalization from the literature in the sense that achievability was shown in literature but exhausting set of all such measurements is not known. We give a method to construct an arbitrary error-minimizing measurement in terms of certain parameters. The method is summarized in a table for the three cases. A quantum example in two parts (see Example 2(a) and 2(b)) is given to illustrate finding error minimizing measurements.
  3. 3.
    For an arbitrary measurement from the set of error minimizing measurements, we have derived the a closed-form expressions of acceptance. We have given a closed-form expression of maximum acceptance obtained from maximizing over set of all error-minimizing measurements in many cases. In some cases, we have given it as an optimization problem.
  4. 4.
    Then we proceed to generalize the results to the case when the support are not equal i.e., $\Pi_{\rho}\neq\Pi_{\sigma}$. In this case, we observe that the minimum-error vanishes. Then, we find the condition on an arbitrary measurement to ensure that the error vanishes. We observe that the error may vanish in three different ways and then write the set of error-minimizing measurements as union of the sets satisfying the three conditions. We discuss the property of the measurements from the three sets. We then shown that two of the sets remain empty, when $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$.
  5. 5.
    We give parameterization for an arbitrary measurements for the three sets. We then find expression of acceptance. In the end, we obtain the expression of maximum acceptance in closed form expression by taking maximum of the maximum acceptance obtained on the three sets.
The notations needed across the paper is given below in this section before describing the system model in the next section.
Notations: Let $\mathcal{H}$ denote a Hilbert’s space. Let $\mathcal{L}(\mathcal{H})$, $\mathcal{R}(\mathcal{H})$, $\mathcal{P}(\mathcal{H})$ and $\mathcal{D}(\mathcal{H})$ respectively denote the set of all linear, hermitian, positive semidefinite and density operator over the Hilbert’s space $\mathcal{H}$. $\mathrm{I}\in\mathcal{P}(\mathcal{H})\subset\mathcal{R}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ is identity transformation. $\nu_{1}-\nu_{2}\in\mathcal{P}(\mathcal{H})$ is also written as $\nu_{1}\geq\nu_{2}$ or $\nu_{2}\leq\nu_{1}$. Any operator $\Pi\in\mathcal{P}(\mathcal{H})$ is a projection operator if $\Pi\Pi=\Pi$. For any $\nu\in\mathcal{R}(\mathcal{H})$, $\Pi_{\nu}$ denotes projection operator onto the eigenspace spanned by set of eigenvectors corresponding to all non-zero eigenvalues of $\nu$. Note that by the properties of projector $\Pi_{\nu}\nu\Pi_{\nu}=\nu$. $\Pi_{\nu}^{\max}$ and $\Pi_{\nu}^{\min}$ denote the projection operator onto eigen-space corresponding to maximum and smallest non-zero eigenvalues of $\nu$, denoted as $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$ respectively. For some projection operator $\Pi\in\mathcal{P}(\mathcal{H})$, $\mathcal{P}(\Pi)\subset\mathcal{P}(\mathcal{H})$ and $\mathcal{S}(\Pi)\subset\mathcal{D}(\mathcal{H})$ denote set of all positive operators and density operators respectively that are invariant w.r.t. $\Pi$ i.e. $\mathcal{P}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{P}(\mathcal{H})\}$ and $\mathcal{S}(\Pi)\stackrel{{\scriptstyle\Delta}}{{=}}\{\nu:\Pi\nu\Pi=\nu,\nu\in\mathcal{D}(\mathcal{H})\}$. For any $\nu\in\mathcal{P}(\mathcal{H})$, $\nu^{-1}$ is an operator obtained by substituting all non-zero eigenvalues of $\nu$ by their inverse. Note that, with this definition of generalized inverse, we get $\nu^{-1}\nu=\Pi_{\nu}$. Given a pair of operators $\nu_{1},\nu_{2}\in\mathcal{P}(\mathcal{H})$, we denote $R_{\max}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty}$ and $R_{\min}(\nu_{1},\nu_{2})=\|\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}\|_{\infty,0}$. This is related to the max relative entropy of the states, which is defined as log of maximum eigenvalue of $\nu_{2}^{-1/2}\nu_{1}\nu_{2}^{-1/2}$ [26].
The paper is organized as follows. Section II describes the system model, the problem statement of testing the two states $\rho$ and $\sigma$, and a derivation of the minimum possible postselected error from the literature before defining the set of error-minimizing measurements and the maximum acceptance over this set. Section III and Section IV contain result for postselected symmetric hypothesis testing problem for the case when $\Pi_{\rho}=\Pi_{\sigma}$ and $\Pi_{\rho}\neq\Pi_{\sigma}$ respectively. In Section III, we begin with deriving the condition on a measurement to achieve the minimum postselected error. Then we characterize the exact set of measurements that achieve minimum postselected error and provide construction for an arbitrary measurement from this set. For better exposition, we give an example to show how the value of acceptance varies over this set and another example to illustrate finding error minimizing measurements. Then, we have maximized acceptance and presented a measurement that achieves the maximizes acceptance. Section IV follow a similar pattern. To maintain the flow, we have just given the key results in the main text while supporting mathematical results are stated and derived in the appendix. Now, we proceed towards describing the system model.
II System Model
We consider a quantum object which can be in one of the two possible states $\rho,\sigma\in\mathcal{D}(\mathcal{H})$. The objective is to determine its state which is unknown. The quantum object being in the state $\rho$ corresponds to null hypothesis and the quantum object being in the state $\sigma$ corresponds to the alternative hypothesis. A measurement is performed with positive operator-valued measure (POVM) $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-(\Lambda_{\rho}+\Lambda_{\sigma})\}$. Note that $\Lambda_{\rho},\Lambda_{\sigma}$ must satisfy the condition $\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})$ and $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ for $\Lambda$ to be a valid measurement. Hence, the set of all such measurements is given by
=$\mathcal{M}\stackrel{{\scriptstyle\Delta}}{{=}}\left\{\Lambda:\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\},\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (1)
On measurement, one of the three outcomes corresponding to the operators $\Lambda_{\rho}$, $\Lambda_{\sigma}$ or $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$ is obtained. The first outcome (corresponding to the operator $\Lambda_{\rho}$) corresponds to accepting the null hypothesis and declaring the unknown state as $\rho$, the second outcome (corresponding to the operator $\Lambda_{\sigma}$) corresponds to accepting the alternative hypothesis and declaring as $\sigma$. If the third outcome (corresponding to the operator $\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}$) occurs, both the hypotheses are rejected. The probabilities of the outcomes corresponding to the object being in the state $\rho$ and $\sigma$ is summarized in the table below.
Unknown state Probability of…
declaring $\rho$ declaring $\sigma$ rejection accepting null or alternative
hypothesis i.e. $\rho$ or $\sigma$
$\rho$ $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ $\mathrm{Tr}\left((\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma})\rho\right)$ $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$
Description — Correct outcome Type-1 error Rejection Acceptance
$\sigma$ $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$ $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ $\mathrm{Tr}\left((\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma})\sigma\right)$ $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$
Description — Type-2 error Correct outcome Rejection Acceptance
If the unknown state is $\rho$ and declared state is $\sigma$, the error is type-1 error and if the unknown state is $\sigma$ and declared state is $\rho$, the error is type- 2 error as mentioned in the table. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$ (collectively denoted as $p$) such that $p_{\rho}>0,p_{\sigma}>0,p_{\rho}+p_{\sigma}=1$, the total error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$.
We define acceptance as the probability of accepting alternative or null hypothesis, i.e. probability of not getting the third outcome. Hence, the acceptance for the state $\rho$ is
$A_{\rho}(\Lambda)\stackrel{{\scriptstyle\Delta}}{{=}}\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})\rho)$ (2)
and for the state $\sigma$, it is
$A_{\sigma}(\Lambda)\stackrel{{\scriptstyle\Delta}}{{=}}\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})\sigma).$ (3)
The definition of acceptance in (2) and (3) gives a feeling that the higher the acceptance, the better the measurement is, as it corresponds to the probability of accepting at least one hypothesis. But, observe that the expression of acceptance consists of two terms: the probability of correct estimation and the error. Intuitively, we desire the probability of correct estimation to be as big as possible and the error as low as possible, hence a trade-off. Hence, our objective is to maximize acceptance only after minimizing the error, i.e., to find the measurement having the highest acceptance over all the measurements that have minimum error.
The postselected probability of an event is defined as the probability of the event conditioned on the event that alternative or null hypothesis is accepted. Given prior probabilities $p_{\rho}$ and $p_{\sigma}$, the error probability is $p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)$ and the probability that the null or alternative hypothesis is accepted is $p_{\rho}A_{\rho}(\Lambda)+p_{\sigma}A_{\sigma}(\Lambda)$. So, the postselected symmetric error $e(\Lambda)$ is defined as
$e(\Lambda)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{p_{\rho}\mathrm{Tr}(\Lambda_{\sigma}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)}{p_{\rho}A_{\rho}(\Lambda)+p_{\sigma}A_{\sigma}(\Lambda)}=\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}.$ (4)
With a little abuse of notation, we have extended definition of $e(\Lambda)$ for any $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma}\}$ such that $\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})$. We denote the set of all such operators as
$\mathcal{O}\stackrel{{\scriptstyle\Delta}}{{=}}\left\{\Lambda:\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma}\},\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})\right\}.$ (5)
Further $\Lambda\in\mathcal{M}$ is called as a measurement or POVM and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ even if it is not mentioned. Similarly, $\Lambda\in\mathcal{O}$ is called as an operator and taken as $\Lambda=\{\Lambda_{\rho},\Lambda_{\sigma}\}$. We now define the minimum postselected symmetric error.
Definition 1 (Minimum postselected symmetric error).
Given prior probability $p_{\rho}$ and $p_{\sigma}$, minimum postselected symmetric error is defined as the minimum achievable postselected symmetric error probability over all measurements i.e.,
$e_{s}(\rho,\sigma,p)\stackrel{{\scriptstyle\Delta}}{{=}}\inf_{\Lambda\in\mathcal{M}}e(\Lambda).$
Recall that $\mathcal{M}$ is the set of all measurements. It is simplified as[25, Theorem 6]
$e_{s}(\rho,\sigma,p)$ $=\inf_{\Lambda\in\mathcal{M}}e(\Lambda)=\inf_{\Lambda\in\mathcal{M}}\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}$
$=\left(1+\sup_{\Lambda\in\mathcal{M}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}\right)^{-1}$
$=\left(1+\sup_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}\right)^{-1}.$ (6)
Now, from [25, Equation (66)], we know that
$\sup_{\Lambda\in\mathcal{O}}{\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}}=\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right).$ (7)
Here, the function $\Xi\left(\nu_{1},\nu_{2}\right)$ is known as Thompson metric[27] and is defined as
$\Xi\left(\nu_{1},\nu_{2}\right)=\begin{cases}\max\{R_{\max}\left(\nu_{1},\nu_{2}\right),R_{\max}\left(\nu_{2},\nu_{1}\right)\},&\text{ if $\Pi_{\nu_{1}}=\Pi_{\nu_{2}}$},\\ \infty,&\text{ otherwise.}\end{cases}$ (8)
Substituting from (7) in (6), the minimum postselected symmetric error is given as [25]
$e_{s}(\rho,\sigma,p)$ $=\left(\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)+1\right)^{-1}.$ (9)
Although the minimum error is given in [25], it is not known how to design measurements to achieve this. In this work, one of the key goals is to characterize an arbitrary measurements, which if performed, obtains the minimum error given by (9). The set of all such error-minimizing measurements is formally defined below.
Figure 2: An illustration showing various sets of measurements. $\mathcal{M}$ is set of all measurements. $\mathcal{E}_{s}(\rho,\sigma,p)$ is set of all error-minimizing measurements. Maximum acceptance measurements comprise the set that achieve maximum acceptance over $\mathcal{E}_{s}(\rho,\sigma,p)$.
Definition 2 (Error-minimizing measurements).
The set of all error minimizing measurements is defined as
$\mathcal{E}_{s}(\rho,\sigma,p)\stackrel{{\scriptstyle\Delta}}{{=}}\{\Lambda:e(\Lambda)=e_{s}(\rho,\sigma,p),\Lambda\in\mathcal{M}\}.$
Within the set of error-minimizing measurements, there are measurements where acceptance for the state $\rho$ and $\sigma$ is maximized (see Fig. 2). The maximum acceptance, denoted as $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, is defined below.
Definition 3 (Maximum acceptance).
The maximum acceptance possible over the set $\mathcal{E}_{s}(\rho,\sigma,p)$ of all error-minimizing measurements is defined as
$A_{\rho}^{s}=\max_{\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)}A_{\rho}(\Lambda)\ \text{ and }\ A_{\sigma}^{s}=\max_{\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)}A_{\sigma}(\Lambda)$
for the states $\rho$ and $\sigma$ respectively.
We further notice that $e_{s}(\rho,\sigma,p)$ vanishes for the states $\rho$, $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$. So, we split the analysis into two parts discussed in next two sections. The first part considers the case $\Pi_{\rho}=\Pi_{\sigma}$ and second part covers the remaining cases. The next section describes the results for the first part.
III The possible states $\rho$ and $\sigma$ have the same support i.e. $\Pi_{\rho}=\Pi_{\sigma}$
We begin the first subsection by finding the condition on the measurement operators that must be satisfied for the minimum error to be achieved. Based on these conditions, we present some key novel properties of the error-minimizing measurements. Further, we write the set of error-minimization measurements in a parameterized form and give a method to construct an arbitrary error-minimizing measurement. Then we give an example in the next subsection to show that the acceptance for an arbitrary measurement varies with the parameters, although all of them being the error-minimizing operators, thus showing the need to maximizing the acceptance. In the last subsection, we give the expression for maximum achievable acceptance for an arbitrary error-minimizing measurement. The following subsection begins with the characterization of error-minimizing measurements.
III-A The set of all error minimizing measurements and an arbitrary construction
The following theorem derives the condition on measurement operators to achieve the minimum postselected symmetric error, along with providing a novel proof of minimum postselected symmetric error.
Theorem III.1.
For any $\rho,\sigma\in\mathcal{D}(\mathcal{H})$, postselected symmetric error is lower bounded as
$e(\Lambda)\geq e_{s}(\rho,\sigma,p)\ \forall\ \Lambda\in\mathcal{O}$ (10)
and the equality is obtained iff measurement operators $\{\Lambda_{\rho},\Lambda_{\sigma}\}$ satisfy the condition:
$\begin{cases}\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\hfill\qquad\underline{\mathcal{C}1}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\hfill\qquad\underline{\mathcal{C}2}\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{otherwise},\hfill\qquad\underline{\mathcal{C}3}\end{cases}$ (11)
where $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$.
Proof.
The proof is given in Appendix B. ∎
Recall from (6) that, minimum of $e(\Lambda)$ can be obtained by first finding $\sup_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$. It turns out that
$\max_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}$ $=\max\left(\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma)},\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{{\mathrm{Tr}(p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\rho}\Lambda_{\sigma}\rho)}\right)$ (12)
$=\max(R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma),R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)).$
The first case in (11) is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies the condition that $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}>\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. In this case $\Lambda_{\rho}$ is taken such that $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$ and $\max_{\Lambda\in\mathcal{O}}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho+p_{\sigma}\Lambda_{\sigma}\sigma)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}=\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{{\mathrm{Tr}(p_{\rho}\Lambda_{\rho}\rho)}}{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma)}=R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)$. Similarly, second case is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies the condition that
$\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}<\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$. The third case is corresponding to ${R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$, and it signifies when $\max_{\Lambda_{\rho}\in\mathcal{P}(\mathcal{H})}\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)}=\max_{\Lambda_{\sigma}\in\mathcal{P}(\mathcal{H})}\frac{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)}{p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)}$, either can be chosen. We have to deal with the cases separately and hence need to represent them compactly. We represent them as $\mathcal{C}1,\mathcal{C}2,$ and $\mathcal{C}3$ respectively for the remaining of this section, as mentioned in (11).
Remark 1 (Notes on $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$).
$\mathrm{T}^{\max}$ denotes the projection operator onto the subspace where $\rho$ is largest as comparison to $\sigma$, in the sense that any vector $|\psi\rangle$ in this subspace, we get $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\max}(\rho,\sigma)$, which is the highest possible value it can have. Similarly for the subspace corresponding to the projection operator $\mathrm{T}^{\min}$, this ratio $\frac{\langle\psi|\rho|\psi\rangle}{\langle\psi|\sigma|\psi\rangle}=R_{\min}(\rho,\sigma)$, which is also the minimum possible value it can achieve. In classical scenario, on taking density operators corresponding to probabilities to be discriminated, $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ would correspond to choosing the realization of random variable for which ratio of the two probabilities are maximum and minimum.
Remark 2 (Notes on $\mathcal{P}(\mathrm{T}^{\max})$ and $\mathcal{P}(\mathrm{T}^{\min})$).
Now, $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\max}$. Similarly, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$ means that all the eigenvectors of $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ corresponding to positive eigenvalues lie in the subspace corresponding to the operator $\mathrm{T}^{\min}$. So, Theorem III.1 states that, by restricting the subspace where eigenvector of the operators $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ lie, one can obtain measurements that achieve minimum postselected symmetric error.
Corollary 1.
If $p,\rho,\sigma$ satisfy $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, then for any $\Lambda$ such that $e(\Lambda)=e_{s}(\rho,\sigma,p)$, one of the following two holds:
  1. 1.
    The measurement never detects $\sigma$ that is $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$. The measurement outcome is either $\rho$ or third outcome i.e. rejecting both the hypotheses (see Figure 3(a)).
  2. 2.
    The measurement never detects $\rho$ that is $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0$. The measurement outcome is either $\sigma$ or third outcome i.e. rejecting both the hypotheses (see Figure 3(b)).
So, for $p,\rho,\sigma$ such that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, any measurement that minimizes postselected symmetric error, either never makes a decision in favour of $\rho$ or never makes a decision in favour of $\sigma$. Fig. 3 depicts the two possible measurements in this case.
Proof.
From the Theorem III.1, when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ we get
$e(\Lambda)=e_{s}(\rho,\sigma,p)\Rightarrow\begin{cases}\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0\Rightarrow\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0,\text{ and }\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\Pi_{\rho}\rho\Pi_{\rho}\right)\\ \qquad=\mathrm{Tr}\left(\Lambda_{\sigma}\Pi_{\sigma}\rho\Pi_{\sigma}\right)=\mathrm{Tr}\left(\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\sigma^{-1/2}\rho\sigma^{-1/2}\right)=0,\\ \text{ or }\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0\Rightarrow\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,\text{ and }\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=\mathrm{Tr}\left(\Lambda_{\rho}\Pi_{\rho}\rho\Pi_{\rho}\right)\\ \qquad=\mathrm{Tr}\left(\Lambda_{\rho}\Pi_{\sigma}\rho\Pi_{\sigma}\right)=\mathrm{Tr}\left(\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\sigma^{-1/2}\rho\sigma^{-1/2}\right)=0.\end{cases}$ (13)
So, for any error-minimizing measurement, one of the two cases must hold. ∎
Figure 3: The figure shows the two possible outcomes when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)\neq R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ for any error minimizing measurement where (a) $\sigma$ never declared, (b) $\rho$ never declared.
Remark 3.
The statement in the Corollary 1 says that the measurements that achieve minimum postselected error, are not detecting one of the states. The observation puts a very fundamental question,“Does a lower postselected error really indicate better hypothesis testing, when prior probability is known?" The observation suggests a negative answer. So, the next question is, “What is a better metric to assess the quality of hypothesis testing, when prior probability is known?" We won’t address it in this work and leave it as an open question.
Corollary 2.
A measurement with non-zero probabilities of detecting both $\rho$ and $\sigma$ is possible only if the prior probabilities $p_{\rho}$ and $p_{\sigma}$ are equal to
$p_{\rho}^{*}=\frac{\sqrt{R_{\max}(\sigma,\rho)}}{\sqrt{R_{\max}(\rho,\sigma)}+\sqrt{R_{\max}(\sigma,\rho)}}\text{ and }p_{\sigma}^{*}=\frac{\sqrt{R_{\max}(\rho,\sigma)}}{\sqrt{R_{\max}(\rho,\sigma)}+\sqrt{R_{\max}(\sigma,\rho)}}.$ (14)
Proof.
From (11), we know that such a measurement is possible when
$R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)$ $=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ (15)
$\Leftrightarrow\frac{p_{\rho}}{p_{\sigma}}R_{\max}(\rho,\sigma)=\frac{p_{\sigma}}{p_{\rho}}R_{\max}(\sigma,\rho)$ (16)
$\Leftrightarrow\frac{p_{\rho}}{p_{\sigma}}=\frac{\sqrt{R_{\max}(\sigma,\rho)}}{\sqrt{R_{\max}(\rho,\sigma)}}.$ (17)
Now, putting the condition that $p_{\rho}+p_{\sigma}=1$, we obtain the desired result. ∎
(a) $p_{\rho}>p_{\rho}^{*}$ and $p_{\sigma}<p_{\sigma}^{*}$
(b) $p_{\rho}<p_{\rho}^{*}$ and $p_{\sigma}>p_{\sigma}^{*}$
(c) $p_{\rho}=p_{\rho}^{*}$ and $p_{\sigma}=p_{\sigma}^{*}$
Figure 4: An illustration showing possible outcomes for different values of prior probability $p_{\rho}$ and $p_{\sigma}$.
From Theorem III.1, we now get three scenarios:
  • $p_{\rho}>p_{\rho}^{*}$ (i.e. $p_{\sigma}<p_{\sigma}^{*}$): Error-minimizing measurement never detects $\sigma$. It either detect $\rho$ or rejects both hypotheses. See Fig. 4(a).
  • $p_{\rho}<p_{\rho}^{*}$ (i.e. $p_{\sigma}>p_{\sigma}^{*}$): Error-minimizing measurement never detects $\rho$. It either detect $\sigma$ or rejects both hypotheses. See Fig. 4(b).
  • $p_{\rho}=p_{\rho}^{*}$ (i.e. $p_{\sigma}=p_{\sigma}^{*}$): There are error-minimizing measurements detecting either/both the hypotheses. See Fig. 4(c).
The observation says that if the prior probability of state being $\rho$ is high, $\sigma$ is never detected under the error minimizing measurement. Similarly, if the prior probability of state being $\sigma$ is high, $\rho$ is never detected. This states that an error minimizing measurement will detect the high probability state or declare nothing. Also, there exists a certain value of the pair $(p_{\rho},p_{\sigma})$ such that an error minimizing measurement can detect both states.
Now, we give two examples. In the first classical example, we show how the previous theorem can be used to find the error minimizing measurements. We also observe that there are measurements, despite having the minimum postselected symmetric error, rejecting both the hypotheses with probability close to $1$. This shows a need to find measurements that maximize acceptance.
In the first example, we take Hilbert’s space with basis given by $\{|0\rangle,|1\rangle,|2\rangle\}$. In this example, we get an intuitive idea when both state $\rho$ and $\sigma$, being a mixture of 3 orthogonal pure states $\{|0\rangle,|1\rangle,|2\rangle\}$. The first state $|0\rangle$ has higher probability in the state $\rho$ as compared to in the state $\sigma$, second orthogonal state $|1\rangle$ has higher probability in the state $\sigma$ as compared to state $\rho$ and third orthogonal state $|2\rangle$ having equal probability in both the states $\rho$ and $\sigma$. In relative terms, observe that, $|0\rangle$ is relatively prominent in the state $\rho$, $|1\rangle$ being prominent in the state $\sigma$, while $|2\rangle$ having equal probability of both. We obtain that the measurements that minimizes error, declares the unknown state as $\rho$ by measuring $|0\rangle\langle 0|$ and $\sigma$ is never detected. We will observe that, for the error-minimizing measurements, the value of acceptances for states $\rho$ and $\sigma$ vary depending on the choice of measurement. For an error-minimizing measurement, the acceptance can be very small, thus not declaring any state with probability close to 1. We suggest that, while designing an error-minimizing measurement, maximizing acceptance should also be considered.
Example 1.
Let $\{|0\rangle,|1\rangle,|2\rangle\}$ be the basis of Hilbert’s space. Take $p_{\rho}=1/2,p_{\sigma}=1/2$,
$\rho=\frac{\mu}{2}|0\rangle\langle 0|+\frac{\mu}{2}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|\text{ and }\sigma=\frac{\mu}{4}|0\rangle\langle 0|+\frac{3\mu}{4}|1\rangle\langle 1|+(1-\mu)|2\rangle\langle 2|.$
Note that, here $\Pi_{\rho}=\Pi_{\sigma}=|0\rangle\langle 0|+|1\rangle\langle 1|+|2\rangle\langle 2|=\mathrm{I}$, so $\mathrm{I}-\Pi_{\sigma}=0$. Now,
$\sigma^{-1/2}\rho\sigma^{-1/2}=2|0\rangle\langle 0|+\frac{2}{3}|1\rangle\langle 1|+|2\rangle\langle 2|\text{ and }\rho^{-1/2}\sigma\rho^{-1/2}=\frac{1}{2}|0\rangle\langle 0|+\frac{3}{2}|1\rangle\langle 1|+|2\rangle\langle 2|.$ (18)
Observe that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(\rho,\sigma)=2$, $R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)=R_{\max}(\sigma,\rho)=3/2$,
$\mathrm{T}^{\max}=\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}=|0\rangle\langle 0|.$
Thus minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=1/3$. Note that $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)>R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, hence, it corresponds to the case $\mathcal{C}1$. We get
$\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=c|0\rangle\langle 0|,\text{ and }\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0.$
So $\Lambda_{\rho}=c\frac{|0\rangle\langle 0|}{\mu/4}=\frac{4c}{\mu}|0\rangle\langle 0|$, $\Lambda_{\sigma}=0$. So, the set of POVM characterizing the measurements that achieve minimum postselected symmetric error is given by
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}=\frac{4c}{\mu}|0\rangle\langle 0|,\Lambda_{\sigma}=0,c\leq\frac{\mu}{4}\right\}$
=$=\left\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:0<c\leq 1,\Lambda_{\rho}=c|0\rangle\langle 0|,\Lambda_{\sigma}=0\right\}.$ (19)
So, an arbitrary error-minimizing measurement can be parameterized as $\{c|0\rangle\langle 0|,0,\mathrm{I}-c|0\rangle\langle 0|\}$ for some $0<c\leq 1$. For these measurements, $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=c\frac{\mu}{2},\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=c\frac{\mu}{4},\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$.
Note that, with this POVM, a conclusive decision is being made for only $c\mu/2$ and $c\mu/4$ fraction of cases when the unknown state is $\rho$ and $\sigma$ respectively. If $\mu$ is small, performance of postselected test would be the same i.e. $e_{s}(\rho,\sigma,p)=1/3$, however we will be declaring inconclusive results in most of the cases, so the measurement would not be really useful. We need some metric to characterize the usefulness of postselected measurement. Second problem is choice of the constant term $c$. We use the metric acceptance to characterize the usefulness.
Now, we give second example. This quantum example is focused towards determining constraints and illustrating various claims made in this subsection about error minimizing measurements.
Example 2(a).
Consider the Hilbert’s space $\mathcal{H}_{2}$ with the basis $\{|0\rangle,|1\rangle\}$. Take the states as
$\rho=\frac{3}{4}|+\rangle\langle+|+\frac{1}{4}|-\rangle\langle-|=\frac{1}{2}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 0|+\frac{1}{4}|0\rangle\langle 1|+\frac{1}{2}|1\rangle\langle 1|,\qquad\sigma=\frac{3}{4}|0\rangle\langle 0|+\frac{1}{4}|1\rangle\langle 1|.$
Note that, in this example $\Pi_{\rho}=\Pi_{\sigma}=\mathrm{I}$. We obtain
$\sigma^{-1/2}\rho\sigma^{-1/2}=\frac{2}{3}|0\rangle\langle 0|+\frac{1}{\sqrt{3}}|1\rangle\langle 0|+\frac{1}{\sqrt{3}}|0\rangle\langle 1|+2|1\rangle\langle 1|.$
We get eigenvalues of $\sigma^{-1/2}\rho\sigma^{-1/2}$ as $\frac{4\pm\sqrt{7}}{3}$ and corresponding eigenvectors as $\frac{{\sqrt{3}|0\rangle+(2\pm\sqrt{7})|1\rangle}}{\sqrt{14\pm 4\sqrt{7}}}.$ Thus $R_{\max}(\rho,\sigma)=\frac{4+\sqrt{7}}{3},R_{\min}(\rho,\sigma)=\frac{4-\sqrt{7}}{3},R_{\max}(\sigma,\rho)=\frac{4+\sqrt{7}}{3}.$ So, we get
$p_{\rho}^{*}=\frac{\sqrt{R_{\max}(\sigma,\rho)}}{\sqrt{R_{\max}(\rho,\sigma)}+\sqrt{R_{\max}(\sigma,\rho)}}=\frac{1}{2},\text{ and similarly }p_{\sigma}^{*}=\frac{1}{2}.$
The projection operators onto the eigenspace corresponding to the maximum and minimum eigenvalue are respectively given by
$\mathrm{T}^{\max}$ $=\frac{1}{14+4\sqrt{7}}\left(\sqrt{3}|0\rangle+(2+\sqrt{7})|1\rangle\right)\left(\sqrt{3}\langle 0|+(2+\sqrt{7})\langle 1|\right)\text{ and }$
$\mathrm{T}^{\min}$ $=\frac{1}{14-4\sqrt{7}}\left(\sqrt{3}|0\rangle+(2-\sqrt{7})|1\rangle\right)\left(\sqrt{3}\langle 0|+(2-\sqrt{7})\langle 1|\right).$
As we obtained $p_{\rho}^{*}=1/2$ and $p_{\sigma}^{*}=1/2$, using Theorem III.1, we obtain minimum error and error minimizing condition as the following.
  1. 1.
    If $p_{\rho}>1/2:e_{s}(\rho,\sigma,p)=\left(1+\frac{p_{\rho}}{p_{\sigma}}\frac{4+\sqrt{7}}{3}\right)^{-1}$ and error minimizing measurement should satisfy the constraint that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$.
  2. 2.
    If $p_{\rho}<1/2:e_{s}(\rho,\sigma,p)=\left(1+\frac{p_{\sigma}}{p_{\rho}}\frac{4+\sqrt{7}}{3}\right)^{-1}$ and error minimizing measurement should satisfy the constraint that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0$, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$.
  3. 3.
    If $p_{\rho}=1/2:e_{s}(\rho,\sigma,p)=\left(1+\frac{4+\sqrt{7}}{3}\right)^{-1}$ and error minimizing measurement should satisfy the constraint that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$, $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$.
Note that $\mathrm{T}^{\max}$ and $\mathrm{T}^{\min}$ are both rank 1 and so, any element in the set $\mathcal{P}(\mathrm{T}^{\max})$ and $\mathcal{P}(\mathrm{T}^{\max})$ are of the form $c\mathrm{T}^{\max}$ and $c\mathrm{T}^{\min}$ for some $c\geq 0$. Following the fact that $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ have simple form, and that $\Pi_{\sigma}=\mathrm{I}$ in this example, we can obtain constraint on $\Lambda_{\rho},\Lambda_{\sigma}$ by simply multiplying $\sigma^{-1/2}$ on both sides, a more general way to find most general form of $\Lambda_{\rho},\Lambda_{\sigma}$ for error-minimization is given next.
Theorem III.1 gives the subspace where $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ lie for $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$ to be an error minimizing measurement. Building on Theorem III.1, we now focus on deriving the set of measurements that achieve minimum postselected symmetric error. We use Lemma C.1 from Appendix C to obtain the constraint on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$, from constraint on $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$ as given in Theorem III.1 to ensure that the measurement achieves the minimum postselected asymmetric error. Then, we parameterize $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in a way to satisfy the obtained constraints and further put conditions $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ to ensure that it remains a valid measurement, thus obtaining the set of all measurements. The error minimizing set and relevant proof is stated in the following theorem.
Theorem III.2.
The set $\mathcal{E}_{s}(\rho,\sigma,p)$ of all measurements achieving the minimum postselected symmetric error for the three cases as mentioned in (11), is given respectively by
  1. $\underline{\mathcal{C}\mathrm{1}}$:
    $\mathcal{E}_{s}(\rho,\sigma,p)=\Bigg{\{}\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\\ \text{ }\hskip 199.16928pt\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
  2. $\underline{\mathcal{C}\mathrm{2}}$:
    $\mathcal{E}_{s}(\rho,\sigma,p)=\Bigg{\{}\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\\ \text{ }\hskip 199.16928pt\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
  3. $\underline{\mathcal{C}\mathrm{3}}$:
    $\mathcal{E}_{s}(\rho,\sigma,p)=\Bigg{\{}\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),c_{r}\in[0,1],\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},c\leq\left\|c_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
Here $\mathrm{P}^{\max}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\max}\sigma^{-1/2}}$ and $\mathrm{P}^{\min}=\Pi_{\sigma^{-1/2}\mathrm{T}^{\min}\sigma^{-1/2}}$ with $\mathrm{T}^{\max}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\max}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ and $\mathrm{T}^{\min}\stackrel{{\scriptstyle\Delta}}{{=}}\Pi^{\min}_{\sigma^{-1/2}\rho\sigma^{-1/2}}$ as defined in Theorem III.1.
Proof.
Using Theorem C.1 and Lemma C.1 from Appendix C, we get
$\begin{gathered}\sigma^{1/2}\Gamma\sigma^{1/2}\in\mathcal{P}(\Pi)\Leftrightarrow\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\Pi_{\sigma^{-1/2}\Pi\sigma^{-1/2}})\text{ and }\\ \sigma^{1/2}\Gamma\sigma^{1/2}=0\Leftrightarrow\Pi_{\sigma}\Gamma\Pi_{\sigma}=0\Leftrightarrow\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}).\end{gathered}$ (22)
We will use these two results for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$ to get the expression for $\Lambda_{\rho}$ and $\Lambda_{\sigma}$. We begin with condition on $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ derived in Theorem III.1 for the three cases separately.
  1. $\underline{\mathcal{C}\mathrm{1}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$. Now, using equivalence shown in (22), we get
    $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\ \Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}).$
    A general $\Lambda_{\rho}$ can be chosen as $\Lambda_{\rho}={c}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}$ for some $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$, and for any such $\Lambda_{\sigma}$ to exist, it has to be ensured that $\Lambda_{\rho}\leq\mathrm{I}$ and thus $c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these condition together, we get an error minimizing measurement as
    $\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}.$
  2. $\underline{\mathcal{C}\mathrm{2}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
    $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}).$
    A general $\Lambda_{\sigma}$ can be chosen as $\Lambda_{\sigma}={c}\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$ for some $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),c\geq 0$. On putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we obtain $\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$, and for any such $\Lambda_{\rho}$ to exist, it has to be ensured that $\Lambda_{\sigma}\leq\mathrm{I}$ and thus $c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\leq\mathrm{I}\Rightarrow c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$. Writing these condition together, we get an error minimizing measurement as
    $\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}.$
  3. $\underline{\mathcal{C}\mathrm{3}}$
    In this case, using Theorem III.1, for error minimizing POVM, we get the condition $\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ and $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min})$. Now, using equivalence shown in (22), we get
    $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}).$
    Here, a general $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ can be chosen as $\Lambda_{\rho}={cc_{r}}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}$ and $\Lambda_{\sigma}={c(1-c_{r})}\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$ for some $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min}),c\geq 0,c_{r}\in[0,1]$ and putting the condition that $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we get
    $\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\displaystyle c\leq\left\|c_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}.$
    Writing this as set, we obtain the result stated in the theorem.
Constructing an arbitrary error minimizing measurement: Theorem III.2 gives an intuitive way to construct any POVM that minimizes error. The way to choose parameters for three cases are considered separately as given in the table below. The first column shows the parameters and rest three show the way to choose them in different cases.
Cases $\mathcal{C}1$ $\mathcal{C}2$ $\mathcal{C}3$
$\psi_{\max}$, $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$, $\psi_{\max}$ not needed, $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),$
$\psi_{\min}$ $\psi_{\min}$ not needed $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})$ $\psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})$
$c_{r}$, $c_{r}$ not needed, $c_{r}$ not needed, $c_{r}\in[0,1]$,
$c$ $c\leq\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}$ $c\leq\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$ $c\leq\left\|\frac{c_{r}\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+\frac{(1-c_{r})\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$
$\Lambda_{\rho},\Lambda_{\sigma}$ $\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},$ $\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},$ $\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},$
$\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$ $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$ $\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$
TABLE I: Table showing a parameterization of an arbitrary error minimizing measurement
The measurement POVM is given as $\{\Lambda_{\sigma},\Lambda_{\rho},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. This is the most general way to construct an arbitrary POVM achieving the minimum postselected symmetric error. This way of constructing the set also gives us more freedom in choosing the measurement in the sense that any possible set of measurements can be obtained by choosing appropriate set for the free variables $\psi_{\max},\psi_{\min},c$ and $c_{r}$.
Remark 4.
For unique characterization of an arbitrary POVM in the set $\mathcal{E}_{s}(\rho,\sigma,p)$, we have:
  1. For the case $\underline{\mathcal{C}\mathrm{1}}$, the set of parameters
    $(c,\psi_{\max},\Lambda_{\sigma})$ is sufficient.
  2. For the case $\underline{\mathcal{C}\mathrm{2}}$, the set of parameters
    $(c,\psi_{\min},\Lambda_{\rho})$ is sufficient.
  3. For the case $\underline{\mathcal{C}\mathrm{3}}$, the set of parameters
    $(c,c_{r},\psi_{\max},\psi_{\min})$ is sufficient.
The above set of values give a parameterization of an arbitrary error minimizing measurement. We will utilize this later when we calculate acceptance in terms of these parameters in a form which can be easily maximized. Interestingly, as we will see, the acceptance turns out to be only a function of $c$ and $c_{r}$.
Finding set of error minimizing measurements for system given in the Example 2(a) is illustrated below. In this example, we will see that maximum value of $c$ can be obtained easily, which can be later utilized to find the maximum acceptance.
Example 2(b).
Referring to system described in Example 2(a), we get
$\mathrm{P}^{\max}$ $=\frac{1}{12+4\sqrt{7}}\left(|0\rangle+(2+\sqrt{7})|1\rangle\right)\left(\langle 0|+(2+\sqrt{7})\langle 1|\right)\text{ and }$
$\mathrm{P}^{\min}$ $=\frac{1}{12-4\sqrt{7}}\left(|0\rangle+(2-\sqrt{7})|1\rangle\right)\left(\langle 0|+(2-\sqrt{7})\langle 1|\right).$
For error minimizing measurement, we have to take $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$ and $\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})$. Note that, $\Pi_{\sigma}=\mathrm{I}$ and $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1, and so we get $\psi_{\max}=\mathrm{P}^{\max}$ and $\psi_{\min}=\mathrm{P}^{\min}$. Further $\mathrm{Tr}\left(\psi_{\max}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right)=\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}$ and $\mathrm{Tr}\left(\psi_{\min}\sigma\right)=\mathrm{Tr}\left(\mathrm{P}_{\min}\sigma\right)=\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}$. So, most general error minimizing measurement is given as below.
  1. 1.
    If $p_{\rho}>1/2:\left\{\{c\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max},0,\mathrm{I}-c\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}\}:0<c\leq\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}\right\}$.
  2. 2.
    If $p_{\rho}<1/2:\left\{\{0,c\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min},\mathrm{I}-c\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\}:0<c\leq\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}\right\}$.
  3. 3.
    If $p_{\rho}=1/2:\\ \left\{\{cc_{r}c\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max},c(1-c_{r})c\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min},\mathrm{I}-cc_{r}\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}-c(1-c_{r})\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\}:\right.\\ \left.\hskip 142.26378pt0\leq c_{r}\leq 1,c\leq\left\|c_{r}\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}+(1-c_{r})\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\right\|_{\infty}^{-1}\right\}$.
In the above description of the set, it is clear that:
  1. 1.
    If $p_{\rho}>1/2:$ maximum value of c is $\frac{1}{4}\frac{14+4\sqrt{7}}{12+4\sqrt{7}}=0.272$.
  2. 2.
    If $p_{\rho}<1/2:$ maximum value of c is $\frac{1}{4}\frac{14-4\sqrt{7}}{12-4\sqrt{7}}=0.603$.
  3. 3.
    If $p_{\rho}=1/2:$ maximum value of c is $\max_{0\leq c_{r}\leq 1}\left\|c_{r}\frac{4(12+4\sqrt{7})}{14+4\sqrt{7}}\mathrm{P}^{\max}-(1-c_{r})\frac{4(12-4\sqrt{7})}{14-4\sqrt{7}}\mathrm{P}^{\min}\right\|_{\infty}^{-1}$.
Here, we are obtaining a simple numerical bound on $c$ in first two cases because $\mathrm{P}^{\max}$ and $\mathrm{P}^{\min}$ are rank 1 operators. Usually, it depends on $\psi_{\max}$ and $\psi_{\min}$. We will see in the next subsection that acceptance depends only on $c$ except for the third case. So, easily obtained maximum value of $c$ helps in finding maximum acceptance.
III-B The maximum acceptance for error-minimizing measurements
Note that Theorem III.2 gives the set of measurements that minimize the postselected symmetric error as $\mathcal{E}_{s}(\rho,\sigma,p)$ and gives their parameterization in terms of $(c,\psi_{\max},\Lambda_{\sigma})$, $(c,\psi_{\min},\Lambda_{\rho})$ or $(c,c_{r},\psi_{\max},\psi_{\min})$. The next Lemma gives acceptance for a given error minimizing measurement as a function of parameters $c$ and $c_{r}$.
Theorem III.3.
For the most general $\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)$, as given in Theorem III.2, the acceptances when the unknown states $\rho$ or $\sigma$, are respectively given below.
  1. For case $\underline{\mathcal{C}\mathrm{1}}$,
    $A_{\rho}(\Lambda)=cR_{\max}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c$.
  2. For case $\underline{\mathcal{C}\mathrm{2}}$,
    $A_{\rho}(\Lambda)=cR_{\min}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c$.
  3. For case $\underline{\mathcal{C}\mathrm{3}}$,
    $A_{\rho}(\Lambda)=cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c$.
Here $c$ and $c_{r}$ are as given in the Table I for an arbitrary error-minimizing measurement.
Proof.
Using Theorem C.1 from Appendix C, we have
$\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})\Rightarrow\sigma^{1/2}\psi_{\max}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}).$
Thus $\mathrm{Tr}\left(\psi_{\max}\rho\right)=\mathrm{Tr}\left(\psi_{\max}\Pi_{\rho}\rho\Pi_{\rho}\right)\stackrel{{\scriptstyle(a)}}{{=}}\mathrm{Tr}\left(\psi_{\max}\Pi_{\sigma}\rho\Pi_{\sigma}\right)$
=$\stackrel{{\scriptstyle(b)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
=$\stackrel{{\scriptstyle(c)}}{{=}}\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}\right)$
=$\stackrel{{\scriptstyle(d)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\mathrm{T}^{\max}\right)$
=$\stackrel{{\scriptstyle(e)}}{{=}}R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\sigma^{1/2}\psi_{\max}\sigma^{1/2}\right).$
Here, $(a)$ uses the condition that $\Pi_{\sigma}=\Pi_{\rho}$. $(b)$ is obtained from $\Pi_{\sigma}=\sigma^{-1/2}\sigma^{1/2}$ and using cyclic property of trace. $(c)$ and $(e)$ uses that $\sigma^{1/2}\psi_{\max}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$. $(d)$ follows from the fact that $\mathrm{T}^{\max}$ is projection onto the subspace corresponding to maximum eigenvalue of $\sigma^{-1/2}\rho\sigma^{-1/2}$ and so, we have $\mathrm{T}^{\max}\sigma^{-1/2}\rho\sigma^{-1/2}=R_{\max}(\rho,\sigma)\mathrm{T}^{\max}$. Thus, we obtain
$\mathrm{Tr}\left(\psi_{\max}\rho\right)=R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\psi_{\max}\sigma\right).$ (23)
In a similar fashion, we can show that
$\mathrm{Tr}\left(\psi_{\min}\rho\right)=R_{\min}(\rho,\sigma)\mathrm{Tr}\left(\psi_{\min}\sigma\right).$ (24)
So, for any $\Gamma$ such that
$\Gamma=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=c\frac{\mathrm{Tr}\left(\psi_{\max}\rho\right)}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}=cR_{\max}(\rho,\sigma)\text{ and }\mathrm{Tr}\left(\Gamma\sigma\right)=c\frac{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}=c.$ (25)
Similarly, for any $\Gamma=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}$, we get
$\mathrm{Tr}\left(\Gamma\rho\right)=cR_{\min}(\rho,\sigma)\text{ and }\mathrm{Tr}\left(\Gamma\sigma\right)=c.$ (26)
On taking
$\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=0\text{ and }\mathrm{Tr}\left(\Gamma\sigma\right)=0.$ (27)
We will use these results from (25),(26), and (27) directly for $\Gamma=\Lambda_{\rho}$ and $\Gamma=\Lambda_{\sigma}$. The set of measurements for which error is minimized are given differently for the three cases, so we derive expression of acceptance for the three cases one-by-one as given below.
  1. $\underline{\mathcal{C}\mathrm{1}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=c\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$ and corresponding acceptance is
    $A_{\rho}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=cR_{\max}(\rho,\sigma)+0=cR_{\max}(\rho,\sigma),$
    $A_{\sigma}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=c+0=c.$
  2. $\underline{\mathcal{C}\mathrm{2}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\sigma}=c\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$ and corresponding acceptance is
    $A_{\rho}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=cR_{\min}(\rho,\sigma)+0=cR_{\min}(\rho,\sigma),$
    $A_{\sigma}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=c+0=c.$
  3. $\underline{\mathcal{C}\mathrm{3}}$
    $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ are parameterized as $\Lambda_{\rho}=cc_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)},\Lambda_{\sigma}=c(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)},\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ and corresponding acceptance is given as
    $A_{\rho}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma),$
    $A_{\sigma}(\Lambda)$ $=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)+\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=c.$
Remark 5.
Barring the case $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, the expression for acceptance only depends on $c$ linearly. In turn, $c$ is upper bounded by a function of $\psi_{\max},\psi_{\min}$ and $c_{r}$. So, for maximizing acceptance, we only need to focus on maximizing $c$ over the set of all error minimizing measurements.
For these identified POVMs that achieve the minimum postselected symmetric error, the maximum acceptance is given in the following theorem.
Theorem III.4.
Under the constraint $e(\Lambda)=e_{s}(\rho,\sigma,p),\Lambda\in\mathcal{M}$, the maximum acceptance for the states $\rho$ and $\sigma$ is given below.
  1. $\underline{\mathcal{C}\mathrm{1}}$
    $A^{s}_{\rho}=R_{\max}(\rho,\sigma)\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right)\quad\text{ and}\quad A^{s}_{\sigma}=\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right).$
  2. $\underline{\mathcal{C}\mathrm{2}}$
    $A^{s}_{\rho}=R_{\min}(\rho,\sigma)\mathrm{Tr}\left(\mathrm{P}^{\min}\sigma\right)\quad\text{ and}\quad A^{s}_{\sigma}=\mathrm{Tr}\left(\mathrm{P}^{\min}\sigma\right).$
  3. $\underline{\mathcal{C}\mathrm{3}}$
    $A^{s}_{\sigma}=\max_{c_{r}\in[0,1]}\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma,c_{r})$ and
    $A^{s}_{\rho}=\max_{c_{r}\in[0,1]}(c_{r}R_{\max}(\rho,\sigma)+(1-c_{r})R_{\min}(\rho,\sigma))\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2},c_{r}).$
Here, for any $\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0$, we have
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}.$
More detailed properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$ are given in Appendix E, which would help finding closed form expression in specific cases.
Proof.
We derive for the three cases separately.
  1. $\underline{\mathcal{C}\mathrm{1}}$
    From Lemma III.3, $A^{s}_{\rho}=R_{\max}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c$ is obtained from Lemma D.2 in Appendix D as stated below
    $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c=\max_{\psi_{\max}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max})}\left\|\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}\right\|_{\infty}^{-1}=\mathrm{Tr}\left(\mathrm{P}^{\max}\sigma\right).$
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, we get the stated results.
  2. $\underline{\mathcal{C}\mathrm{2}}$
    From Lemma III.3, $A^{s}_{\rho}=R_{\min}(\rho,\sigma)\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c\qquad\text{ and }\qquad A^{s}_{\sigma}=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c.$ Now $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c$ is obtained from Lemma D.2 in Appendix D as stated below
    $\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c=\max_{\psi_{\min}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})}\left\|\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}=\mathrm{Tr}\left(\mathrm{P}^{\min}\sigma\right).$
    Substituting it in the expression of $A_{\rho}^{s}$ and $A_{\sigma}^{s}$, we get the stated results.
  3. $\underline{\mathcal{C}\mathrm{3}}$
    From Lemma III.3, $A_{\rho}(\Lambda)=cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma)\text{ and }A_{\sigma}(\Lambda)=c.$ Note that these functions have to be maximized with respect to $c$ and $c_{r}$. Taking first and maximizing $c$ we get
    $A^{s}_{\rho}$ $=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}cc_{r}R_{\max}(\rho,\sigma)+c(1-c_{r})R_{\min}(\rho,\sigma)$
    $=\max_{c_{r}\in[0,1]}c_{r}R_{\max}(\rho,\sigma)+(1-c_{r})R_{\min}(\rho,\sigma)\max_{\begin{subarray}{c}\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\\ \psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})\end{subarray}}\left\|\frac{{c_{r}}\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+\frac{\left(1-c_{r}\right)\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$
    $=\max_{c_{r}\in[0,1]}(c_{r}R_{\max}(\rho,\sigma)+(1-c_{r})R_{\min}(\rho,\sigma))\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma,c_{r}).$
    Here the last step is obtained from Lemma D.3 in Appendix D. Now maximizing $A_{\sigma}(\Lambda)$, we obtain
    $A_{\sigma}^{s}$ $=\max_{\mathcal{E}_{s}(\rho,\sigma,p)}c=\max_{c_{r}\in[0,1]}\max_{\begin{subarray}{c}\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\\ \psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})\end{subarray}}\left\|c_{r}\frac{\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+(1-c_{r})\frac{\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}^{-1}$
    $=\max_{c_{r}\in[0,1]}\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2},c_{r}).$
    Here the last step is obtained from Lemma D.3 in Appendix D.
Remark 6.
Note that for the case $\mathcal{C}1$, the expressions for acceptance for $\rho$ and $\sigma$, both are linear in the parameter $c$ in Theorem III.3, so the parameters that maximize acceptance for both states $\rho$ and $\sigma$ is the same. Thus, same measurement maximizes both the acceptances. This also holds for the case $\mathcal{C}2$. In contrast, in the case $\mathcal{C}3$, we obtained acceptances as different functions of $c$ and $c_{r}$. So, the measurements which maximize acceptance for the state $\rho,i.e.A_{\rho}(\Lambda)$, won’t be maximizing acceptance for the state $\sigma,i.e.A_{\sigma}(\Lambda)$ and the measurement which maximize the acceptance for the state $\sigma,i.e.A_{\sigma}(\Lambda)$ won’t be maximizing acceptance for the state $\rho,i.e.A_{\rho}(\Lambda)$.
Remark 7.
The maximum acceptance can be obtained by taking the measurement described by the POVM $\Lambda=\{\mathrm{P}^{\max},0,\mathrm{I}-\mathrm{P}^{\max}\}$ and $\Lambda=\{0,\mathrm{P}^{\min},\mathrm{I}-\mathrm{P}^{\min}\}$ in $\mathcal{C}1$ and $\mathcal{C}2$ respectively. In $\mathcal{C}3$, finding the measurement, which achieves the maximum acceptance for the state $\rho$ (or $\sigma$) is involved. First step is finding $c_{r}$ such that $A_{\rho}^{s}$ (or $A_{\sigma}^{s}$) is obtained and placing corresponding optimal $\psi_{\max}$ and $\psi_{\min}$ and $c=\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}(\sigma,c_{r})$, maximum acceptance achieving measurement can be obtained. This requires solving the optimization step and finding the solution.
This completes deriving the expression for maximum acceptance over the set of measurements which achieves minimum postselected symmetric error. In the next section, we maximize the acceptance for the pair of states $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$.
IV The possible states $\rho$ and $\sigma$ do not have the same support i.e. $\Pi_{\rho}\neq\Pi_{\sigma}$
Recall from (9) that the minimum postselected symmetric error $e_{s}(\rho,\sigma,p)=\left(\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)+1\right)^{-1}$. If $\Pi_{\rho}\neq\Pi_{\sigma}$, from the definition given in (8), it turns out that $\Xi\left(p_{\rho}\rho,p_{\sigma}\sigma\right)=\infty$, hence $e_{s}(\rho,\sigma,p)=0$. In this section, we begin with finding the condition on $\Lambda$ to obtain $e(\Lambda)=e_{s}(\rho,\sigma,p)=0$. The following theorem states the set of all possible error-minimizing measurements as a union of three sets, with each corresponding to one of the three conditions needed to ensure that the error vanishes. An arbitrary measurement must belong to one of three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$ defined in the Theorem below to ensure that $e(\Lambda)=0$.
Theorem IV.1.
For any measurement $\Lambda\in\mathcal{M}$, $e(\Lambda)=0$ if and only if
$\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)\cup\mathcal{E}^{2}_{s}(\rho,\sigma)\cup\mathcal{E}^{3}_{s}(\rho,\sigma),$ (28)
where $\mathcal{E}^{1}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$, $\mathcal{E}^{2}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$, $\mathcal{E}^{3}_{s}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\{\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$.
Proof.
The proof is given in Appendix F. ∎
Remark 8.
The three sets described in the Theorem IV.1 have a specific structure of outcomes, which is described below (also see Fig. 5).
  1. 1.
    The first set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ i.e. probability of detecting $\rho$ when the given state is $\rho$ is non-vanishing and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\rho$, the measurement either declares in favor of $\rho$ or rejects both the hypotheses. However, if the given state is $\sigma$, it always rejects both the hypotheses. Fig. 5(a) illustrates this point.
  2. 2.
    Similarly, the second set i.e. $\mathcal{E}_{s}^{2}(\rho,\sigma)$ consists of all the measurements for which only $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ i.e. probability of detecting $\sigma$ when the given state is $\sigma$ is not vanishing and the probabilities $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$, $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)$, and $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ vanish. So, for any measurement that belongs to this set, if the given input state is $\sigma$, the measurement either declares in favor of $\sigma$ or rejects both the hypotheses. However, if the given state is $\rho$, it always rejects both the hypotheses. Fig. 5(b) illustrates this point.
  3. 3.
    The third set i.e. $\mathcal{E}_{s}^{3}(\rho,\sigma)$ consists of all the measurements for which both $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ are non-zero. So, for the given state $\rho$ (or $\sigma$), it either declares outcome as $\rho$ (or $\sigma$), or rejects both the hypotheses. Fig. 5(c) illustrates this point.
(a) $\mathcal{E}^{1}_{s}(\rho,\sigma)$
(b) $\mathcal{E}^{2}_{s}(\rho,\sigma)$
(c) $\mathcal{E}^{3}_{s}(\rho,\sigma)$
Figure 5: The figure shows possible outcomes for an arbitrary measurement from the set $\mathcal{E}^{1}_{s}(\rho,\sigma)$, $\mathcal{E}^{2}_{s}(\rho,\sigma)$ and $\mathcal{E}^{3}_{s}(\rho,\sigma)$.
Remark 9.
Note that for any measurement for which error vanishes, $\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0$ or equivalently $\mathrm{Tr}(\Lambda_{\rho}\sigma)=0$ and $\mathrm{Tr}(\Lambda_{\sigma}\rho)=0$ and so both type-1 and type-2 error has to vanish. This follows from the fact that numerator of $e(\Lambda)$ has to vanishes to ensure that error vanishes. So, it either declares the correct outcome or reject both the hypotheses but never declares the wrong outcome. Further note that, one of $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ i.e. at least one of the probability of correct decision has to be non-zero to ensure that denominator in $e(\Lambda)$ remains non-zero.
Remark 10.
Note that the set and respective conditions depend only the subspace spanned by eigenvectors of $\rho$ and $\sigma$ and not on the prior probabilities $p_{\rho}$ and $p_{\sigma}$. So, error-minimizing measurements are independent of the prior probabilities when $\Pi_{\rho}\neq\Pi_{\sigma}$.
Corollary 3.
If $\Pi_{\sigma}<\Pi_{\rho},\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. So, the set of all error minimizing measurement is given only by first set i.e. $\mathcal{E}_{s}^{1}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\sigma$, $A_{\sigma}(\Lambda)=0$, so if the input state is $\sigma$, the error-minimizing measurement always rejects both the hypotheses. Also, any error minimizing measurement never declares the unknown state as $\sigma$. See Fig. 5(a) for clarity.
Proof.
If $\Pi_{\sigma}\leq\Pi_{\rho}\Rightarrow\Pi_{\rho+\sigma}=\Pi_{\rho}$. So, if $\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),$ then $\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ and so $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$, so $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ are empty sets and
$\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}_{s}^{1}(\rho,\sigma).$
Now, note that for any $\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0$ and $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0$, so $A_{\sigma}(\Lambda)=0$ for any error minimizing measurement. ∎
Corollary 4.
If $\Pi_{\rho}<\Pi_{\sigma},\mathcal{E}_{s}(\rho,\sigma)=\mathcal{E}^{2}_{s}(\rho,\sigma)$ and $A_{\rho}(\Lambda)=0$. So, the set of all error minimizing measurement is given only by the second set i.e. $\mathcal{E}_{s}^{2}(\rho,\sigma)$. Further, for the postselected symmetric error minimizing measurements, acceptance for the state $\rho$, $A_{\rho}(\Lambda)=0$, so if the input state is $\rho$, the error-minimizing measurement always rejects both the hypotheses. Also, any error minimizing measurement never declares the unknown state as $\rho$. See Fig. 5(b) for clarity.
Proof.
Follows similar to the proof of Corollary 3. ∎
We now give a parameterization of the three sets in the following three lemmas.
Lemma IV.1.
The set $\mathcal{E}^{1}_{s}(\rho,\sigma)$ is completely characterized as $\mathcal{E}^{1}_{s}(\rho,\sigma)=$
$\Bigg{\{}\Lambda:$ $\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),0<c_{1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
Proof.
Starting from the definition of $\mathcal{E}_{s}^{1}(\rho,\sigma)$ as given in Theorem IV.1, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$
$\Leftrightarrow\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},c_{1}>0,\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\rho}\leq\mathrm{I},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$
$\Leftrightarrow\displaystyle\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),0<c_{1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1},\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho},\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
Writing it as a set, we get the stated result in the Lemma. ∎
Lemma IV.2.
The set $\mathcal{E}^{2}_{s}(\rho,\sigma)$ is completely characterized as $\mathcal{E}^{2}_{s}(\rho,\sigma)=$
$\Bigg{\{}\Lambda$ $:\Lambda_{\sigma}=\frac{c_{2}\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),0<c_{2}\leq\left\|\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
Proof.
Starting from the definition of $\mathcal{E}_{s}^{2}(\rho,\sigma)$ as given in Theorem IV.1, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$
$\Leftrightarrow\displaystyle\Lambda_{\sigma}=\frac{c_{2}\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},c_{2}>0,\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),\Lambda_{\sigma}\leq\mathrm{I},\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$
$\Leftrightarrow\displaystyle\Lambda_{\sigma}=\frac{c_{2}\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),0<c_{2}\leq\left\|\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1},\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma},\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}).$
Writing it as a set, we get the stated result in the Lemma. ∎
Lemma IV.3.
The set $\mathcal{E}^{3}_{s}(\rho,\sigma)$ is completely characterized as
$\mathcal{E}^{3}_{s}(\rho,\sigma)=\Bigg{\{}\Lambda:\Lambda_{\rho}=\frac{c_{3}c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),$ $\ \Lambda_{\sigma}=\frac{c_{3}(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),$
$c_{r}\in[0,1],0<c_{3}\leq\left\|\frac{c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\Bigg{\}}.$
Proof.
Starting from the definition of $\mathcal{E}_{s}^{3}(\rho,\sigma)$ as given in Theorem IV.1, we get
$\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}.$
Taking $\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho}),c_{r}\in[0,1],c_{3}>0$, the above mentioned conditions can be parameterized as
$\Lambda_{\rho}=\frac{c_{3}c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\Lambda_{\sigma}=\frac{c_{3}(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}.$
Now substituting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the condition $\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$, we get this condition as a bound on $c_{3}$ as
$\Lambda_{\rho}=\frac{c_{3}c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\Lambda_{\sigma}=\frac{c_{3}(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},c_{3}\leq\left\|\frac{c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}.$
Writing it as a set, we get the stated result in the Lemma. ∎
Remark 11.
Note that all the three cases, taking $\psi_{\rho}$ and $\psi_{\sigma}$ such that $\mathrm{Tr}\left(\psi_{\rho}\rho\right)=0$ or $\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)=0$ is not valid, as clear from the expression of $\Lambda_{\rho}$ and $\Lambda_{\sigma}$. This ensures that $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$. This is needed because $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0$ or $\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$ is among the conditions on the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ in Theorem IV.1.
Constructing an arbitrary error minimizing measurement: In Theorem IV.1, we saw that an arbitrary error-minimizing measurement $\Lambda$ must belong to one of $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$. The outcome specific properties of these sets were observed in Remark 8 and illustrated in Fig. 5. Further, the sets were written in a parameterized form in the Lemma IV.1, Lemma IV.2 and Lemma IV.3. Based on these parameters, a method to construct an arbitrary error minimizing measurement is given in the table below. The first column in the table states the free variables/parameters to choose and rest give the condition on these. The conditions for $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$ are written in second, third and fourth columns respectively. To construct an arbitrary error-minimizing measurement from $\mathcal{E}_{s}^{1}(\rho,\sigma)$, start from the first row, pick an arbitrary free variable satisfying the conditions stated the second column and proceed towards the last row in this way. After getting $\Lambda_{\rho}$ and $\Lambda_{\sigma}$ in the last row, the error-minimizing measurement is given by $\{\Lambda_{\rho},\Lambda_{\sigma},\mathrm{I}-\Lambda_{\rho}-\Lambda_{\sigma}\}$. For constructing measurements from the sets $\mathcal{E}_{s}^{2}(\rho,\sigma)$ or $\mathcal{E}_{s}^{3}(\rho,\sigma)$, do the same but take the conditions from the corresponding columns.
Set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ $\mathcal{E}_{s}^{2}(\rho,\sigma)$ $\mathcal{E}_{s}^{3}(\rho,\sigma)$
$\psi_{\rho}$, $\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})$, Not needed, $\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})$,
$\psi_{\sigma}$ Not needed $\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho})$ $\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho})$
$c_{r}$, Not needed, Not needed, $c_{r}\in[0,1]$,
$c_{\{\cdot\}}$ $c_{1}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}$ $c_{2}\leq\left\|\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}$ $c_{3}\leq\left\|\frac{c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}$
$\Lambda_{\rho},\Lambda_{\sigma}$ $\Lambda_{\rho}=c_{1}\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},$ $\Lambda_{\sigma}=c_{2}\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},$ $\Lambda_{\rho}=c_{3}c_{r}\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},$
$\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\leq\mathrm{I}-\Lambda_{\rho}$ $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\rho}\leq\mathrm{I}-\Lambda_{\sigma}$ $\Lambda_{\sigma}=c_{3}(1-c_{r})\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}$
TABLE II: The table shows a way to choose parameters for an arbitrary error-minimizing measurement for each of the three sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$.
Now, we give the acceptance for an arbitrary error minimizing measurement from the three sets in terms of the parameters given in Table II.
Theorem IV.2.
For an arbitrary error minimizing measurement $\Lambda\in\mathcal{E}_{s}(\rho,\sigma)$, acceptance for the states $\rho$ and $\sigma$ is given as
  • If $\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma)$: $A_{\rho}(\Lambda)=c_{1}$ and $A_{\sigma}(\Lambda)=0$.
  • If $\Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)$: $A_{\rho}(\Lambda)=0$ and $A_{\sigma}(\Lambda)=c_{2}$.
  • If $\Lambda\in\mathcal{E}_{s}^{3}(\rho,\sigma)$: $A_{\rho}(\Lambda)=c_{3}c_{r}$ and $A_{\sigma}(\Lambda)=c_{3}(1-c_{r})$.
Here $c_{1},c_{2},c_{3}$ and $c_{r}$ are parameters as given in the Table II.
Proof.
We know that
$\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\rho})\Rightarrow\mathrm{Tr}\left(\Gamma\rho\right)=0\text{ and }\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})\Rightarrow\mathrm{Tr}\left(\Gamma\sigma\right)=0.$
[…]
For these three sets, the acceptance is calculated as given below.
  • If $\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma)$, then $\Lambda_{\rho}=\frac{c_{1}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\psi_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ and so
    $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=c_{1},\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0\quad\Rightarrow A_{\rho}(\Lambda)=c_{1},A_{\sigma}(\Lambda)=0.$
  • If $\Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)$, then $\Lambda_{\sigma}=c_{2}\frac{\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ and so
    $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=c_{2}\quad\Rightarrow A_{\rho}(\Lambda)=0,A_{\sigma}(\Lambda)=c_{2}.$
  • If $\Lambda\in\mathcal{E}_{s}^{3}(\rho,\sigma)$, then $\Lambda_{\rho}=\frac{c_{3}c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)},\Lambda_{\sigma}=\frac{c_{3}(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)},\psi_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\psi_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho})$ and so
    $\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=c_{3}c_{r},\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ $=0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=c_{3}(1-c_{r})$
    $\Rightarrow A_{\rho}(\Lambda)=c_{3}c_{r},A_{\sigma}(\Lambda)=c_{3}(1-c_{r}).$
Now, we maximize the acceptance over the the set $\mathcal{E}\left(\rho,\sigma,p\right)$. To do so, we maximize the expression obtained in Theorem IV.2 with respect the parameters. The maximum acceptance expression is derived in the following theorem.
Theorem IV.3 (Maximum acceptance).
Given a pair of states as $\rho$ and $\sigma$ such that $\Pi_{\rho}\neq\Pi_{\sigma}$, then maximum acceptance for the states $\rho$ and $\sigma$ is given as
$\rho,\sigma$ such that $A_{\rho}^{s}$ $A_{\sigma}^{s}$
$\Pi_{\rho}>\Pi_{\sigma}$ $1-\mathrm{Tr}\left(\Pi_{\sigma}\rho\right)$ 00
$\Pi_{\rho}<\Pi_{\sigma}$ 00 $1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right)$
$\Pi_{\rho}\not<\Pi_{\sigma}$ and $\Pi_{\rho}\not>\Pi_{\sigma}$ $1-\mathrm{Tr}\left(\Pi_{\sigma}\rho\right)$ $1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right)$
Proof.
Recall from Corollary 3 that when $\Pi_{\rho}>\Pi_{\sigma}$, the set $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)$. Similarly, from Corollary 4, we know that when $\Pi_{\rho}<\Pi_{\sigma}$, the set $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{2}_{s}(\rho,\sigma)$. We begin with the third case i.e. $\Pi_{\rho}\not<\Pi_{\sigma}$ and $\Pi_{\rho}\not>\Pi_{\sigma}$. In this case $\mathcal{E}_{s}(\rho,\sigma,p)=\mathcal{E}^{1}_{s}(\rho,\sigma)\cup\mathcal{E}^{2}_{s}(\rho,\sigma)\cup\mathcal{E}^{3}_{s}(\rho,\sigma)$. Starting from maximizing acceptance for the state $\rho$ over the set $\mathcal{E}_{s}(\rho,\sigma,p)$, which can be obtained by maximizing over all the three sets followed by taking the maximum of the three, we get
$A_{\rho}^{s}=\max_{\Lambda\in\mathcal{E}_{s}(\rho,\sigma,p)}A_{\rho}(\Lambda)$ $=\max\left(\max_{\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma)}A_{\rho}(\Lambda),\max_{\Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)}A_{\rho}(\Lambda),\max_{\Lambda\in\mathcal{E}_{s}^{3}(\rho,\sigma)}A_{\rho}(\Lambda)\right).$ (29)
Note that $A_{\rho}(\Lambda)=0\ \forall\ \Lambda\in\mathcal{E}_{s}^{2}(\rho,\sigma)$, hence
$A_{\rho}^{s}$ $=\max\left(\max_{\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma)}A_{\rho}(\Lambda),\max_{\Lambda\in\mathcal{E}_{s}^{3}(\rho,\sigma)}A_{\rho}(\Lambda)\right).$ (30)
Using the expression of acceptance obtained in Theorem IV.2, we get
$A_{\rho}^{s}$ $=\max\left(\max_{\Lambda\in\mathcal{E}_{s}^{1}(\rho,\sigma)}c_{1},\max_{\Lambda\in\mathcal{E}_{s}^{3}(\rho,\sigma)}c_{3}c_{r}\right).$ (31)
Applying bounds on $c_{1}$ and $c_{3}$ in the set $\mathcal{E}_{s}^{1}(\rho,\sigma)$ and $\mathcal{E}_{s}^{3}(\rho,\sigma)$, we get
$A_{\rho}^{s}$ $=\max\left(\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1},\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho})}c_{r}\left\|\frac{c_{r}\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\right).$ (32)
Taking $c_{r}$ inside the max-norm expression
$A_{\rho}^{s}$ $=\max\left(\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1},\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}),\psi_{\sigma}\in\mathcal{S}(\mathrm{I}-\Pi_{\rho})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}^{-1}\right).$ (33)
Note that $\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\leq\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\Rightarrow\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}\leq\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}+\frac{(1-c_{r})\psi_{\sigma}}{c_{r}\mathrm{Tr}\left(\psi_{\sigma}\sigma\right)}\right\|_{\infty}$. Thus, the maximum is obtained by maximizing the first term and so
$A_{\rho}^{s}$ $=\max_{\psi_{\rho}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}=\max_{\psi_{\rho}\in\mathcal{S}(\Pi_{\rho+\sigma}-\Pi_{\sigma})}\left\|\frac{\psi_{\rho}}{\mathrm{Tr}\left(\psi_{\rho}\rho\right)}\right\|_{\infty}^{-1}.$ (34)
The equality can be proved similar to proof of (75) in proof of Lemma D.2 of Appendix D. Note that among the all $\tilde{\psi}_{\rho}\in\mathcal{P}(\Pi_{\rho+\sigma}-\Pi_{\sigma})$, such that $\|\tilde{\psi}_{\rho}\|_{\infty}=1$, the one that maximizes $\mathrm{Tr}\left(\tilde{\psi}_{\rho}\rho\right)$ is $\tilde{\psi}_{\rho}=(\Pi_{\rho+\sigma}-\Pi_{\sigma})$. Hence, we obtain maximum at $\psi_{\rho}=\frac{\Pi_{\rho+\sigma}-\Pi_{\sigma}}{\mathrm{Tr}\left(\Pi_{\rho+\sigma}-\Pi_{\sigma}\right)}$ and is given by
$A_{\rho}^{s}$ $=\mathrm{Tr}\left((\Pi_{\rho+\sigma}-\Pi_{\sigma}))\rho\right)=1-\mathrm{Tr}\left(\Pi_{\sigma}\rho\right).$ (35)
Acceptance for the state $\sigma$ is maximized in a similar way to obtain.
$A_{\sigma}^{s}$ $=1-\mathrm{Tr}\left(\Pi_{\rho}\sigma\right).$ (36)
For the case $\Pi_{\rho}<\Pi_{\sigma}$ or $\Pi_{\rho}>\Pi_{\sigma}$, the maximum acceptance can be obtained in a similar way by maximizing over the sets $\mathcal{E}_{s}^{1}(\rho,\sigma)$ or $\mathcal{E}_{s}^{2}(\rho,\sigma)$ respectively, which make the problem simpler and we directly get expression as in (34), thus obtaining the stated values in the table. ∎
Remark 12.
In all the three cases, the stated maximum acceptance for the state $\rho$, i.e. $A_{\rho}^{s}$ is achieved by taking the measurement $\{\Pi_{\rho+\sigma}-\Pi_{\sigma},0,\mathrm{I}-\Pi_{\sigma}+\Pi_{\rho+\sigma}\}$. In all the three cases, the stated maximum acceptance for the state $\sigma$, i.e. $A_{\sigma}^{s}$ is achieved by taking the measurement $\{0,\Pi_{\rho+\sigma}-\Pi_{\rho},\mathrm{I}-\Pi_{\rho}+\Pi_{\rho+\sigma}\}$.
V Conclusion
In this work, we start with giving the set of operators that minimize the postselected symmetric error. We make an important observation about the error-minimizing measurements that such measurements never decide in favor of one of the possible states unless the prior probability has a particular value given by $(p_{\rho}^{*},p_{\sigma}^{*})$ in (14). For a lower prior probability of either of the states, any error-minimizing measurement never declares that state. This puts a fundamental question on minimizing postselected symmetric error $e(\Lambda)$ as a way to find the best measurement. After this, we have given an arbitrary construction of an error-minimizing measurement in terms of freely chosen variables. We then show by an example that the value of acceptance varies for different measurements taken from the set of error-minimizing measurements, despite all being error-minimizing. This example illustrates the need for maximizing the acceptance over the set of error-minimizing measurements. Further, we have given the expression of acceptance for an arbitrary error-minimizing measurement in terms of free variables used for constructing the error-minimizing measurement. This is followed by stating the expression for maximum acceptance in closed form-expression except for the case when $R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$ by maximizing the expression of acceptance with respect to the free variable. Here, we would also add that, all these results are generalizable in a more simplified form in case of discriminating classical probability distributions on taking appropriate density matrices and measurements.
The maximum acceptance obtained here over the set of error-minimizing measurements means that any lower value of acceptance is achievable with the same minimum error probability, but if higher acceptance is desired, the error probability is bound to increase. Thus, the work opens up the question of characterizing postselected symmetric error under the demand of higher acceptance. This work also opens up other new potential directions for research for example, studying acceptance in the asymptotic case. It may lead to a new class of quantum divergence parameterized by acceptance for any pair of quantum states. Recently postselected communication over quantum channels was proposed in[28] and capacity is derived. Some metric similar to acceptance in this work is potentially needed there as well. Similarly, methods of finding appropriate constrains can be generalized to problems on probabilistic protocols as in [29, 30, 31]. Overall, this work opens up a new arena to explore other matrices for performance of postselected symmetric hypothesis testing problems going beyond error probability and studying acceptance in related problems.
Appendix A Condition for equality
The appendix focuses on deriving certain certain relation between an operator $\nu\in\mathcal{P}(\mathcal{H})$ and projectors $\Pi_{\nu},\Pi_{\nu}^{\max}$, and $\Pi_{\nu}^{\min}$ and eigenvalues $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$. With the help of these properties we have given bounds on $\mathrm{Tr}\left(\zeta\nu\right)$ for $\zeta\in\mathcal{S}(\mathcal{H})$ or with more constraints on $\zeta$. These bounds are derived in Lemma A.3 and Lemma A.6, which will be utilized in deriving bounds in Appendix B in the proof of Theorem III.1.
We take $\nu\in\mathcal{P}(\mathcal{H})$ as fixed operator and $\zeta\in\mathcal{P}(\mathcal{H})$ as a variable operator with more constraints if needed. Eigenvalue decomposition of $\nu$ and projection operator of the subspace spanned by eigenvectors is given as
$\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|\text{ and }\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|,$ (37)
where $k_{\nu,i}$ represents eigenvalue and $|e_{i}\rangle$ is corresponding eigenvector of $\nu$. The projection operators corresponding to the maximum and minimum eigenvalues are given as
$\Pi^{\max}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|\text{ and }\Pi^{\min}_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\text{ respectively.}$. (38)
Recall that $\|\nu\|_{\infty}$ and $\|\nu\|_{\infty,0}$ denote the maximum and minimum eigenvalue of $\nu$ respectively. We will use this notation throughout Appendix A.
Lemma A.1.
For any $\nu\in\mathcal{P}(\mathcal{H})$, the following holds:
  1. $(\mathrm{1})$
    $\Pi_{\nu}^{\max}\Pi_{\nu}=\Pi_{\nu}^{\max}$,
  2. $(\mathrm{2})$
    $\Pi_{\nu}^{\max}\nu=\|\nu\|_{\infty}\Pi_{\nu}^{\max}$,
  3. $(\mathrm{3})$
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=1\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$,
  4. $(\mathrm{4})$
    $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$,
  5. $(\mathrm{5})$
    $\Pi_{\nu}^{\min}\Pi_{\nu}=\Pi_{\nu}^{\min}$,
  6. $(\mathrm{6})$
    $\Pi_{\nu}^{\min}\nu=\|\nu\|_{\infty,0}\Pi_{\nu}^{\min}$,
  7. $(\mathrm{7})$
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=1\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$,
  8. $(\mathrm{8})$
    $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
Proof.
Proof of each statement is given below.
  1. $(\mathrm{1})$
    Note that $\Pi_{\nu}=\sum_{i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}^{\max}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|$. Hence,
    $\Pi_{\nu}^{\max}\Pi_{\nu}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}\sum_{j}|e_{i}\rangle\langle e_{i}||e_{j}\rangle\langle e_{j}|=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|=\Pi_{\nu}^{\max}.$
  2. $(\mathrm{2})$
    $\nu=\sum_{i}k_{\nu,i}|e_{i}\rangle\langle e_{i}|$ and $\Pi_{\nu}^{\max}=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|$. Hence
    $\Pi_{\nu}^{\max}\nu$ =$=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}|e_{i}\rangle\langle e_{i}|k_{\nu,j}|e_{j}\rangle\langle e_{j}|=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\sum_{j}\mathds{1}(j=i)k_{\nu,j}|e_{j}\rangle\langle e_{j}|$
    =$=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}k_{\nu,i}|e_{i}\rangle\langle e_{i}|=\sum_{\{i:k_{\nu,i}=\|\nu\|_{\infty}\}}\|\nu\|_{\infty}|e_{i}\rangle\langle e_{i}|=\|\nu\|_{\infty}\Pi_{\nu}^{\max}.$
  3. $(\mathrm{3})$
    From the definition of set $\mathcal{S}(\Pi_{\nu}^{\max})$, it follows that $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\zeta$ and $\mathrm{Tr}\left(\zeta\right)=1$. So,
    $\mathrm{Tr}\left(\zeta\Pi_{\nu}\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\Pi_{\nu}\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\right)=\mathrm{Tr}\left(\zeta\right)=1.$ (39)
    First, third and fourth equality above follows from the definition of $\mathcal{S}(\Pi_{\nu}^{\max})$ and second equality follows from $(\mathrm{1})$.
  4. $(\mathrm{4})$
    Similar to the proof of $(\mathrm{3})$, using $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\zeta$ and $\mathrm{Tr}\left(\zeta\right)=1$, we get
    $\mathrm{Tr}\left(\zeta\nu\right)=\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\nu\right)=\|\nu\|_{\infty}\mathrm{Tr}\left(\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}\right)=\|\nu\|_{\infty}\mathrm{Tr}\left(\zeta\right)=\|\nu\|_{\infty}.$ (40)
    First, third and fourth equality above follows from the definition of $\mathcal{S}(\Pi_{\nu}^{\max})$ and second equality follows from $(\mathrm{2})$.
  5. $(\mathrm{5})$
    --$(\mathrm{8})$ Proof is similar to proof of $(\mathrm{1})-(\mathrm{4})$.
This completes the proof of all statements mentioned in the Lemma. ∎
Lemma A.2.
Given $\nu\in\mathcal{P}(\mathcal{H}),$$\langle\phi|\nu|\phi\rangle\leq\|\nu\|_{\infty}\ \forall\ |\phi\rangle\in\mathcal{H}$ and $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$.
Proof.
We first prove the inequality and then show the equivalence of equality. Using eigenvalue decomposition of $\nu$,
$\langle\phi|\nu|\phi\rangle=\sum_{i}\langle\phi|k_{\nu,i}|e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\langle\phi||e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}.$ (41)
Note that $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=\langle\phi|\left(\sum_{i}|e_{i}\rangle\langle e_{i}|\right)|\phi\rangle\leq\langle\phi|I|\phi\rangle=\langle\phi|\phi\rangle=1$ and $k_{\nu,i}\leq\|\nu\|_{\infty}\ \forall i$. Hence
$\langle\phi|\nu|\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\leq\sum_{i}\|\nu\|_{\infty}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty}\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\leq\|\nu\|_{\infty}.$ (42)
So $\langle\phi|\nu|\phi\rangle\leq\|\nu\|_{\infty}$, which is the inequality in the lemma.
For equality to hold, both the inequalities in (42) have to be equality. First inequality becomes equality iff $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$ and second inequality becomes equality iff $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1$. So we get
$\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1.$ (43)
Now, $\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=1\Leftrightarrow\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=1\Leftrightarrow\langle\phi|\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle=1\Leftrightarrow\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle=|\phi\rangle.$ So we get
$\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle.$ (44)
On substituting $\lvert\langle e_{i}|\phi\rangle|=0\ \text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$ in $|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle$, we get
$|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}|e_{i}\rangle\langle e_{i}|\phi\rangle=\Pi_{\nu}^{\max}|\phi\rangle.$ (45)
For showing the equivalence, converse is also needed, which is proved below.
$|\phi\rangle=\Pi_{\nu}^{\max}|\phi\rangle=\!\!\!\!\sum_{i:k_{\nu,i}=\|\nu\|_{\infty}}\!\!\!\!|e_{i}\rangle\langle e_{i}|\phi\rangle\Rightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\text{ and }|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle.$ (46)
So, we get $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle\text{ with }\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$. Hence, $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi\rangle=|\phi\rangle$. ∎
Lemma A.3.
For any $\nu\in\mathcal{P}(\mathcal{H})$, $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}\ \forall\ \zeta\in\mathcal{D}(\mathcal{H})$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$.
Proof.
The proof is done in 2 steps. First we have shown that $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$, followed by showing $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}$. Combining these two, we get the stated result.
  1. 1.
    Take eigenvalue decomposition of $\zeta=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ with $\sum_{i}k_{\zeta,i}=1$. Then $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle$. Using Lemma A.2, $\langle\phi_{i}|\nu|\phi_{i}\rangle\leq\|\nu\|_{\infty}$ and $\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty}\Leftrightarrow\Pi_{\nu}^{\max}|\phi_{i}\rangle=|\phi_{i}\rangle$. Hence,
    $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle\leq\|\nu\|_{\infty}\sum_{i}k_{\zeta,i}=\|\nu\|_{\infty},$ (47)
    with $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Leftrightarrow\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty}\ \forall\ i\Leftrightarrow\Pi_{\nu}^{\max}|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall\ i$. Hence,
    $\Pi_{\nu}^{\max}\zeta\Pi_{\nu}^{\max}=\sum_{i}k_{\zeta,i}\Pi_{\nu}^{\max}|\phi_{i}\rangle\langle\phi_{i}|\Pi_{\nu}^{\max}=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|=\zeta\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max}).$ (48)
    So, $\mathrm{Tr}\left(\zeta\nu\right)\leq\|\nu\|_{\infty}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})$.
  2. 2.
    Using statement $(\mathrm{4})$ of Lemma A.1, we get
    $\zeta\in\mathcal{S}(\Pi_{\nu}^{\max})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}.$ (49)
Combining the two, we get the desired result. ∎
Lemma A.4.
$\zeta\in\mathcal{S}(\Pi)\Leftrightarrow\Pi\zeta=\zeta\Leftrightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$, where $\{|\phi_{i}\rangle\}$ is set of eigenvectors of $\zeta$.
Proof.
We first prove the first equivalence using the definition of $\mathcal{S}(\Pi)$.
$\zeta\in\mathcal{S}(\Pi)\Rightarrow\zeta=\Pi\zeta\Pi=\Pi\Pi\zeta\Pi=\Pi\zeta.$ (50)
$\zeta=\Pi\zeta\Rightarrow\zeta=\zeta\Pi\Rightarrow\zeta=\Pi\zeta\Pi\Rightarrow\zeta\in\mathcal{S}(\Pi).$ (51)
This completes the proof of first equivalence. Let $k_{\zeta,i}\neq 0$ denote the eigenvalue corresponding to eigenvector $|\phi_{i}\rangle$, then $\forall i$
$\Pi\zeta=\zeta\Rightarrow\Pi\zeta|\phi_{i}\rangle=\zeta|\phi_{i}\rangle\Rightarrow k_{\zeta,i}\Pi|\phi_{i}\rangle=k_{\zeta,i}|\phi_{i}\rangle\Rightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall\ i.$ (52)
Now taking from the right side,
$\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$ $\Rightarrow\sum_{i}k_{\zeta,i}\Pi|\phi_{i}\rangle\langle\phi_{i}|\Pi=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ (53)
$\Rightarrow\Pi\left(\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|\right)\Pi=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ (54)
$\Rightarrow\Pi\zeta\Pi=\zeta$ (55)
$\Rightarrow\zeta\in\mathcal{S}(\Pi).$ (56)
Thus $\zeta\in\mathcal{S}(\Pi)\Leftrightarrow\Pi\zeta=\zeta\Leftrightarrow\Pi|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall i$. ∎
Lemma A.5.
Given $\nu\in\mathcal{P}(\mathcal{H}),$$\langle\phi|\nu|\phi\rangle\geq\|\nu\|_{\infty,0}\ \forall\ |\phi\rangle:\Pi_{\nu}|\phi\rangle=|\phi\rangle$ and $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty}\text{ at }\Pi_{\nu}|\phi\rangle=|\phi\rangle\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$.
Proof.
We first prove the inequality and then show the equivalence of equality by showing the following 3 conditions as equivalent.
  1. 1.
    $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}$,
  2. 2.
    $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}$,
  3. 3.
    $\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$.
Using eigenvalue composition of $\nu$,
$\langle\phi|\nu|\phi\rangle=\sum_{i}\langle\phi|k_{\nu,i}|e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\langle\phi||e_{i}\rangle\langle e_{i}||\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}.$ (57)
Note that $k_{\nu,i}\geq\|\nu\|_{\infty,0}\ \forall i$ and
$\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\sum_{i}\langle\phi|e_{i}\rangle\langle e_{i}|\phi\rangle=\langle\phi|\Pi_{\nu}|\phi\rangle=\langle\phi|\Pi_{\nu}\Pi_{\nu}^{\min}|\phi\rangle=\langle\phi|\Pi_{\nu}^{\min}|\phi\rangle=\langle\phi\|\phi\rangle=1.$
Hence,
$\langle\phi|\nu|\phi\rangle=\sum_{i}k_{\nu,i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}\geq\sum_{i}\|\nu\|_{\infty,0}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty,0}\sum_{i}\lvert\langle e_{i}|\phi\rangle\rvert^{2}=\|\nu\|_{\infty,0}.$ (58)
This proves the inequality of lemma.
Now equality holds iff $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}$. This proves the equivalence with statement 2. Note that $\Pi_{\nu}|\phi\rangle=|\phi\rangle$ and so $|\phi\rangle=\sum_{i}|e_{i}\rangle\langle e_{i}|\phi\rangle$. On substituting $\lvert\langle e_{i}|\phi\rangle|=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty}$, we get
$|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\phi\rangle=\Pi_{\nu}^{\min}|\phi\rangle.$ (59)
For showing the equivalence, converse is also needed, which is proved below.
$|\phi\rangle=\Pi_{\nu}^{\min}|\phi\rangle=\sum_{i:k_{\nu,i}=\|\nu\|_{\infty,0}}|e_{i}\rangle\langle e_{i}|\phi\rangle\Rightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}.$ (60)
So, we conclude $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}\text{ at }\Pi_{\nu}|\phi\rangle=|\phi\rangle\Leftrightarrow\langle e_{i}|\phi\rangle=0\text{ for any }i:k_{\nu,i}\neq\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$. Hence, $\langle\phi|\nu|\phi\rangle=\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi\rangle=|\phi\rangle$. ∎
Lemma A.6.
For any $\nu\in\mathcal{P}(\mathcal{H})$, $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}\ \forall\ \zeta\in\mathcal{S}(\Pi_{\nu})$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Leftrightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
Proof.
The proof is done in 2 steps. First we have shown that $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$, followed by showing $\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}$. Combining these two, we get the desired result.
  1. 1.
    Take eigenvalue decomposition of $\zeta=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|$ with $\sum_{i}k_{\zeta,i}=1$. Then $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle$. From Lemma A.4, we know that $\Pi_{\nu}|\phi_{i}\rangle=|\phi_{i}\rangle$. Now, using Lemma A.5, $\langle\phi_{i}|\nu|\phi_{i}\rangle\geq\|\nu\|_{\infty,0}$ and $\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty,0}\Leftrightarrow\Pi_{\nu}^{\min}|\phi_{i}\rangle=|\phi_{i}\rangle$. Hence,
    $\mathrm{Tr}\left(\zeta\nu\right)=\sum_{i}k_{\zeta,i}\langle\phi_{i}|\nu|\phi_{i}\rangle\geq\|\nu\|_{\infty,0}\sum_{i}k_{\zeta,i}=\|\nu\|_{\infty},$ (61)
    with $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Leftrightarrow\langle\phi_{i}|\nu|\phi_{i}\rangle=\|\nu\|_{\infty,0}\ \forall\ i\Leftrightarrow\Pi_{\nu}^{\min}|\phi_{i}\rangle=|\phi_{i}\rangle\ \forall\ i$. Hence,
    $\Pi_{\nu}^{\min}\zeta\Pi_{\nu}^{\min}=\sum_{i}k_{\zeta,i}\Pi_{\nu}^{\min}|\phi_{i}\rangle\langle\phi_{i}|\Pi_{\nu}^{\min}=\sum_{i}k_{\zeta,i}|\phi_{i}\rangle\langle\phi_{i}|=\zeta\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min}).$ (62)
    So, $\mathrm{Tr}\left(\zeta\nu\right)\geq\|\nu\|_{\infty,0}$ and $\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}\Rightarrow\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})$.
  2. 2.
    Using statement $(\mathrm{8})$ of Lemma A.1, we get
    $\zeta\in\mathcal{S}(\Pi_{\nu}^{\min})\Rightarrow\mathrm{Tr}\left(\zeta\nu\right)=\|\nu\|_{\infty,0}.$ (63)
Combining the two, we get the desired result. ∎
Appendix B Proof of Theorem III.1
$e(\Lambda)$ can be written as $e(\Lambda)=\left(1+R_{\Lambda,p}(\rho,\sigma)\right)^{-1}$, where
$R_{\Lambda,p}(\rho,\sigma)\stackrel{{\scriptstyle\Delta}}{{=}}\frac{p_{\rho}\mathrm{Tr}(\Lambda_{\rho}\rho)+p_{\sigma}\mathrm{Tr}(\Lambda_{\sigma}\sigma)}{p_{\sigma}\mathrm{Tr}(\Lambda_{\rho}\sigma)+p_{\rho}{\mathrm{Tr}(\Lambda_{\sigma}\rho)}}.$ (64)
On taking $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}$ and $\Lambda_{\sigma}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$,
$\mathrm{Tr}\left(\sigma^{-1/2}\Lambda_{\rho}^{s}\sigma^{-1/2}\rho\right)$ $=\mathrm{Tr}\left(\Pi_{\sigma}\Lambda_{\rho}\Pi_{\sigma}\rho\right)\stackrel{{\scriptstyle(a)}}{{=}}\mathrm{Tr}\left(\Pi_{\rho}\Lambda_{\rho}\Pi_{\rho}\rho\right)=\mathrm{Tr}\left(\Lambda_{\rho}\Pi_{\rho}\rho\Pi_{\rho}\right)=\mathrm{Tr}\left(\Lambda_{\rho}\rho\right),$
where $(a)$ is due to $\Pi_{\rho}=\Pi_{\sigma}$. Also, $\mathrm{Tr}\left(\Lambda_{\rho}^{s}\right)=\mathrm{Tr}(\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2})=\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right).$ Similarly, $\mathrm{Tr}\left(\sigma^{-1/2}\Lambda_{\sigma}^{s}\sigma^{-1/2}\rho\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)$ and $\mathrm{Tr}\left(\Lambda_{\sigma}^{s}\right)=\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$. On substituting these in $R_{\Lambda}(p_{\rho}\rho,p_{\sigma}\sigma)$, we get
$R_{\Lambda,p}$ $(\rho,\sigma)=\frac{p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}^{s}\sigma^{-1/2}\rho\sigma^{-1/2}\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}^{s}\right)}{p_{\sigma}\mathrm{Tr}\left(\Lambda_{\rho}^{s}\right)+p_{\rho}\mathrm{Tr}\left(\Lambda_{\sigma}^{s}\sigma^{-1/2}\rho\sigma^{-1/2}\right)}.$
Choosing $\Lambda_{\rho}^{s}=c_{\rho}\Lambda_{\rho}^{s,c}$ and $\Lambda_{\sigma}^{s}=c_{\sigma}\Lambda_{\sigma}^{s,c}$ such that trace of $\Lambda_{\rho}^{s,c}$ and $\Lambda_{\sigma}^{s,c}$ is $1$. We get
$R_{\Lambda,p}(\rho,\sigma)$ $=\frac{p_{\sigma}c_{\sigma}+p_{\rho}c_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)}{p_{\sigma}c_{\rho}+p_{\rho}c_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)}=\frac{c_{\sigma}+c_{\rho}\frac{p_{\rho}}{p_{\sigma}}\mathrm{Tr}\left(\Lambda_{\rho}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)}{c_{\rho}+c_{\sigma}\frac{p_{\rho}}{p_{\sigma}}\mathrm{Tr}\left(\Lambda_{\sigma}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)}.$ (65)
Using Lemma A.3 and Lemma A.6 from Appendix A, we get
$\mathrm{Tr}\left(\Lambda_{\rho}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)\leq\|\sigma^{-1/2}\rho\sigma^{-1/2}\|_{\infty}\text{ with equality iff }\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})\text{ and}$
$\mathrm{Tr}\left(\Lambda_{\sigma}^{s,c}\sigma^{-1/2}\rho\sigma^{-1/2}\right)\geq\|\sigma^{-1/2}\rho\sigma^{-1/2}\|_{\infty,0}\text{ with equality iff }\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min}).$
Substituting these bounds in the expression of $R_{\Lambda,p}(\rho,\sigma)$ in Equation (65), we get upper bound as
$R_{\Lambda,p}$ $(\rho,\sigma)\leq\frac{c_{\sigma}+c_{\rho}R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}{c_{\rho}+c_{\sigma}R_{\min}(p_{\rho}\rho,p_{\sigma}\sigma)}=R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)\frac{c_{\sigma}+c_{\rho}R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}{c_{\sigma}+c_{\rho}R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},$ (66)
with the equality being obtained iff $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ and $\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min})$. Now, ignoring the constant multiplier $R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)$, RHS of (66) is maximized as
$\max_{c_{\rho},c_{\sigma}}\frac{c_{\sigma}+c_{\rho}R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}{c_{\sigma}+c_{\rho}R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}=\begin{cases}\frac{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ 1,&\text{otherwise}.\end{cases}$ (67)
The maximum in (67) being obtained at $\underset{c_{\rho},c_{\sigma}}{\arg\max}\frac{c_{\sigma}+c_{\rho}R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}{c_{\sigma}+c_{\rho}R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}$
=$=\begin{cases}c_{\rho}>0,c_{\sigma}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}=0,c_{\sigma}>0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ c_{\rho}\geq 0,c_{\sigma}\geq 0,(c_{\rho},c_{\sigma})\neq(0,0),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}={R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}.\end{cases}$ (68)
Combining (66) and (67), we obtain
$R_{\Lambda,p}(\rho,\sigma)$$\leq\begin{cases}{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)},&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ {R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},&\text{otherwise}\end{cases}$ (69)
$=\max({R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)},{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}).$ (70)
So, $\min e(\Lambda)=(1+\max({R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)},{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)}))^{-1}$, which is one of the desired results.
Further, the minimum of $e(\Lambda)$ being achieved at $c_{\rho}$ and $c_{\sigma}$ satisfying (68) and $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ and $\Lambda_{\sigma}^{s,c}\in\mathcal{S}(\mathrm{T}^{\min})$. Combining $\Lambda_{\rho}^{s,c}\in\mathcal{S}(\mathrm{T}^{\max})$ with $c_{r}=0$ and $c_{r}>0$ gives $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0$ and $\Lambda_{\rho}^{s}=\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max})$ respectively. Similarly, we get condition on $\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}$. Thus, the final condition for equality in (69) and so minimum of $e(\Lambda)$ is obtained as
$\begin{cases}\sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}=0,&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}>{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}=0,\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{if }{R_{\max}(p_{\rho}\rho,p_{\sigma}\sigma)}<{R_{\max}(p_{\sigma}\sigma,p_{\rho}\rho)},\\ \sigma^{1/2}\Lambda_{\rho}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\max}),\quad\sigma^{1/2}\Lambda_{\sigma}\sigma^{1/2}\in\mathcal{P}(\mathrm{T}^{\min}),&\text{otherwise},\end{cases}$
which is the desired result.
Appendix C Generalized form of operator given the projector
Lemma C.1.
For any $\zeta\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$, the following three are equivalent:
$(\mathrm{1})\ \Pi\zeta\Pi=0\qquad(\mathrm{2})\ \Pi\zeta=\zeta\Pi=0\qquad(\mathrm{3})\ \zeta\in\mathcal{P}(\mathrm{I}-\Pi).$
Proof.
We will prove $(\mathrm{1})\Rightarrow(\mathrm{2})\Rightarrow(\mathrm{3})\Rightarrow(\mathrm{1})$.
(1$\to$(2): Let $\zeta=\sum_{i}k_{i}|\psi_{i}\rangle\langle\psi_{i}|$ be eigenvalue decomposition of $\zeta$ with $k_{i}>0$, then
$\Pi\zeta\Pi=0\Rightarrow\mathrm{Tr}\left(\Pi\zeta\right)=0\Rightarrow\sum_{i}k_{i}\langle\psi_{i}|\Pi|\psi_{i}\rangle=0.$
Note that $k_{i}>0$ and $\langle\psi_{i}|\Pi|\psi_{i}\rangle\geq 0$, hence $\sum_{i}k_{i}\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$ iff $\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$. So we obtain
$\Pi\zeta\Pi=0\Rightarrow\langle\psi_{i}|\Pi|\psi_{i}\rangle=0\ \forall\ i\stackrel{{\scriptstyle(a)}}{{\Rightarrow}}\Pi|\psi_{i}\rangle=0\ \forall\ i\ \Rightarrow\Pi\zeta=\zeta\Pi=0.$
$(a)$ follows from the fact that, if norm squared of a vector $\Pi|\psi_{i}\rangle$, which is $\langle\psi_{i}|\Pi|\psi_{i}\rangle=0$, then the vector $\Pi|\psi_{i}\rangle=0$.
(2$\to$(3): $\Pi\zeta=\zeta\Pi=0\Rightarrow(\mathrm{I}-\Pi)\zeta=\zeta=\zeta(\mathrm{I}-\Pi)\Rightarrow\zeta=(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi)$.
(3$\to$(1): $\zeta\in\mathcal{P}(\mathrm{I}-\Pi)\Rightarrow\zeta=(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Rightarrow\Pi\zeta\Pi=\Pi(\mathrm{I}-\Pi)\zeta(\mathrm{I}-\Pi)\Pi=(\Pi-\Pi)\zeta(\Pi-\Pi)=0$. So, we get equivalence of three statements. ∎
Lemma C.2.
For any $\zeta\in\mathcal{P}(\mathcal{H})$ and $\nu\in\mathcal{P}(\mathcal{H})$,
$\mathrm{Tr}\left(\zeta\nu\right)=0\Leftrightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}).$
Proof.
Note that $\zeta,\nu\in\mathcal{P}(\mathcal{H})$, hence we have $\nu^{1/2}\zeta\nu^{1/2}\geq 0$. Now
$\mathrm{Tr}\left(\zeta\nu\right)=0\stackrel{{\scriptstyle(a)}}{{\Leftrightarrow}}\mathrm{Tr}\left(\nu^{1/2}\zeta\nu^{1/2}\right)=0\stackrel{{\scriptstyle(b)}}{{\Leftrightarrow}}\nu^{1/2}\zeta\nu^{1/2}=0\stackrel{{\scriptstyle(c)}}{{\Leftrightarrow}}\Pi_{\nu}\zeta\Pi_{\nu}=0\stackrel{{\scriptstyle(d)}}{{\Leftrightarrow}}\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}).$
$(a)$ follows from cyclic property of $\mathrm{Tr}\left(\cdot\right)$ operator. $(b)$ follows from the fact that $\nu^{1/2}\zeta\nu^{1/2}\geq 0$. $(c)$ follows from multiplying both side by $\nu^{-1/2}$ or $\nu^{1/2}$ (for reverse direction). $(d)$ utilizes Lemma C.1. ∎
Lemma C.3.
For any $\nu\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$ such that $\Pi\nu=\nu$, then
$\zeta\in\mathcal{P}(\Pi_{\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi).$
Proof.
By taking hermitian and multiplying with $\nu^{-1}$ and $\nu$, we obtain
$\Pi\nu=\nu\Leftrightarrow\nu\Pi=\nu\Leftrightarrow\Pi\Pi_{\nu}=\Pi_{\nu}\Leftrightarrow\Pi_{\nu}\Pi=\Pi_{\nu}.$
Now, beginning with $\zeta\in\mathcal{P}(\Pi_{\nu})$, we get
$\zeta=\Pi_{\nu}\zeta\Pi_{\nu}=\Pi\Pi_{\nu}\zeta\Pi_{\nu}\Pi=\Pi\zeta\Pi\Rightarrow\zeta\in\mathcal{P}(\Pi).$
Hence, $\zeta\in\mathcal{P}(\Pi_{\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi).$ This completes the proof. ∎
Lemma C.4.
For any $\nu\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$ such that $\Pi\Pi_{\nu}=\Pi_{\nu}\Pi=\Pi$, then
$\mathrm{P}\Pi_{\nu}=\mathrm{P},\text{ where }\mathrm{P}=\Pi_{\nu^{-1}\Pi\nu^{-1}}.$
Proof.
Using the definition of projector, we obtained that $\mathrm{P}=\Pi_{\nu^{-1}\Pi\nu^{-1}}=\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\left(\nu^{-1}\Pi\nu^{-1}\right)$ and $\nu^{-1}\Pi_{\nu}=\nu^{-1}$. Now beginning wit $\mathrm{P}\Pi_{\nu}$, we get
$\mathrm{P}\Pi_{\nu}$ $=\Pi_{\nu^{-1}\Pi\nu^{-1}}\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}(\nu^{-1}\Pi\nu^{-1})\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}\Pi_{\nu}$
$=(\nu^{-1}\Pi\nu^{-1})^{-1}\nu^{-1}\Pi\nu^{-1}=\mathrm{P},$
which is the desired result. ∎
Lemma C.5.
For any $\nu\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$ such that $\Pi\Pi_{\nu}=\Pi_{\nu}\Pi=\Pi$, then
$\zeta\in\mathcal{P}(\Pi_{\nu\mathrm{P}\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi),$
where $\mathrm{P}=\Pi_{\nu^{-1}\Pi\nu^{-1}}=\left(\nu^{-1}\Pi\nu^{-1}\right)\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}$.
Proof.
Note that $\nu\mathrm{P}\nu=\nu\left(\nu^{-1}\Pi\nu^{-1}\right)\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\nu={\Pi_{\nu}\Pi\nu^{-1}}\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\nu$, hence
$\Pi\nu\mathrm{P}\nu={\Pi\Pi_{\nu}\Pi\nu^{-1}}\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\nu={\Pi_{\nu}\Pi\nu^{-1}}\left(\nu^{-1}\Pi\nu^{-1}\right)^{-1}\nu=\nu\mathrm{P}\nu.$
So we have $\Pi\nu\mathrm{P}\nu=\nu\mathrm{P}\nu$. Now, using the Lemma C.3, we obtain $\zeta\in\mathcal{P}(\Pi_{\nu\mathrm{P}\nu})\Rightarrow\zeta\in\mathcal{P}(\Pi).$
Lemma C.6.
For projector $\mathrm{P}$ such that $\mathrm{P}\Pi_{\nu}=\mathrm{P}$, then
$\zeta\in\mathcal{P}(\mathrm{P})\Rightarrow\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}}).$
Proof.
Note that $\zeta\in\mathcal{P}(\mathrm{P})$ and $\mathrm{P}\Pi_{\nu}=\mathrm{P}$, hence
$(\nu^{-1}\mathrm{P}\nu^{-1})(\nu\zeta\nu)(\nu^{-1}\mathrm{P}\nu^{-1})=\nu^{-1}\mathrm{P}\Pi_{\nu}\zeta\Pi_{\nu}\mathrm{P}\nu^{-1}=\nu^{-1}\mathrm{P}\zeta\mathrm{P}\nu^{-1}=(\nu^{-1}\zeta\nu^{-1}).$
Now, to prove $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$,
$\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}}(\nu^{-1}\zeta\nu^{-1})\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}}$ $=\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}}(\nu^{-1}\mathrm{P}\nu^{-1})(\nu\zeta\nu)(\nu^{-1}\mathrm{P}\nu^{-1})\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}(\nu^{-1}\mathrm{P}\nu^{-1})\nu\zeta\nu(\nu^{-1}\mathrm{P}\nu^{-1})=\nu^{-1}\zeta\nu^{-1}.$
$(a)$ is obtained from the fact that $\eta\Pi_{\eta}=\Pi_{\eta}\eta=\eta$ on taking $\eta=\nu^{-1}\mathrm{P}\nu^{-1}$. Thus we get $\nu^{-1}\zeta\nu^{-1}\in\mathcal{P}(\Pi_{\nu^{-1}\mathrm{P}\nu^{-1}})$, which completes the proof. ∎
Theorem C.1.
Given $\Pi,\nu\in\mathcal{P}(\mathcal{H}),\Pi\Pi_{\nu}=\Pi$, then the following statements are equivalent
  1. $(\mathrm{1})$
    $\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi)$,
  2. $(\mathrm{2})$
    $\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})$,
  3. $(\mathrm{3})$
    $\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\nu}+\mathrm{P})$,
where $\mathrm{P}=\Pi_{\nu^{-1/2}\Pi\nu^{-1/2}}$.
Proof.
We will first prove $(1)\Leftrightarrow(2)$ and then $(2)\Leftrightarrow(3)$.
(1$\to$(2): Can be obtained just by appropriate substitution in Lemma C.6.
(2$\to$(1): Using Lemma C.6, we get $\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})\Rightarrow\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi_{\nu^{1/2}\mathrm{P}\nu^{1/2}}).$ Now using Lemma C.5, we get $\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi_{\nu^{1/2}\mathrm{P}\nu^{1/2}})\Rightarrow\nu^{1/2}\zeta\nu^{1/2}\in\mathcal{P}(\Pi).$
(2$\to$(3): Using Lemma C.4, we know that $\Pi\Pi_{\nu}=\Pi\Rightarrow\mathrm{P}=\mathrm{P}\Pi_{\nu}$. Hence
$\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})$ $\Rightarrow\mathrm{P}\Pi_{\nu}\zeta\Pi_{\nu}\mathrm{P}=\mathrm{P}\Pi_{\nu}\zeta\Pi_{\nu}=\Pi_{\nu}\zeta\Pi_{\nu}\mathrm{P}=\Pi_{\nu}\zeta\Pi_{\nu}$
$\Rightarrow\mathrm{P}\zeta\mathrm{P}=\Pi_{\nu}\zeta\mathrm{P}=\mathrm{P}\zeta\Pi_{\nu}=\Pi_{\nu}\zeta\Pi_{\nu}$
$\Rightarrow(\Pi_{\nu}-\mathrm{P})\zeta(\Pi_{\nu}-\mathrm{P})=0.$
Using Lemma C.1, $(\Pi_{\nu}-\mathrm{P})\zeta(\Pi_{\nu}-\mathrm{P})=0\Rightarrow(\Pi_{\nu}-\mathrm{P})\zeta=\zeta(\Pi_{\nu}-\mathrm{P})=0$ and so
$\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{P}(\mathrm{P})$ $\Rightarrow(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta(\mathrm{I}-\Pi_{\nu}+\mathrm{P})=\zeta\Rightarrow\zeta\in\mathcal{S}(\mathrm{I}-\Pi_{\nu}+\mathrm{P}).$
(3$\to$(2): Using Lemma C.4, we know that $\Pi\Pi_{\nu}=\Pi\Rightarrow\mathrm{P}=\mathrm{P}\Pi_{\nu}$. Hence
$\zeta\in\mathcal{S}(\mathrm{I}-\Pi_{\nu}+\mathrm{P})$ $\Rightarrow\zeta=(\mathrm{I}-\Pi_{\nu}+\mathrm{P})\zeta$
$\Rightarrow(\Pi_{\nu}-\mathrm{P})\zeta=\zeta(\Pi_{\nu}-\mathrm{P})=0$
$\Rightarrow\Pi_{\nu}\zeta=\mathrm{P}\zeta,\zeta\Pi_{\nu}=\zeta\mathrm{P}$
$\Rightarrow\mathrm{P}\Pi_{\nu}\zeta\Pi_{\nu}\mathrm{P}=\mathrm{P}\zeta\mathrm{P}=\Pi_{\nu}\zeta\Pi_{\nu}$
$\Rightarrow\Pi_{\nu}\zeta\Pi_{\nu}\in\mathcal{S}(\mathrm{P}).$
which completes the proof. ∎
Theorem C.2.
For any $\sigma\in\mathcal{P}(\mathcal{H})$ and projector $\mathrm{P}$ such that $\mathrm{P}\Pi_{\sigma}=\mathrm{P}$ then,
  1. $(\mathrm{1})$
    $\zeta\in\mathcal{P}(\mathrm{P})\Rightarrow\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$ and so $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$.
  2. $(\mathrm{2})$
    $\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\Rightarrow\Pi_{\sigma}\zeta\Pi_{\sigma}\in\mathcal{P}(\mathrm{P})$ and so $\{\Pi_{\sigma}\zeta\Pi_{\sigma}:\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}\subseteq\mathcal{P}(\mathrm{P})$.
Proof.
We will prove $\mathrm{(1)}$ and then $\mathrm{(2)}$.
[…]
(2): Beginning with the definition of $\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$ and simplifying further, we get
$\zeta\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\Rightarrow\zeta=(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\zeta\Rightarrow\Pi_{\sigma}\zeta=\mathrm{P}\zeta=\mathrm{P}\Pi_{\sigma}\zeta.$
[…]
Appendix D Minimizing the max norm
Lemma D.1.
For any $\nu\in\mathcal{P}(\mathcal{H})$ and projector $\Pi$, $\|\Pi\nu\Pi\|_{\infty}\leq\|\nu\|_{\infty}$.
Proof.
Note that $\langle\phi|\Pi\nu\Pi|\phi\rangle=\langle\phi|\Pi|\phi\rangle\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\leq\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}.$ Beginning with the definition of max-norm and using this inequality, we get
$\|\Pi\nu\Pi\|_{\infty}$ $=\max_{|\phi\rangle}\langle\phi|\Pi\nu\Pi|\phi\rangle\leq\max_{|\phi\rangle}\frac{\langle\phi|\Pi}{\sqrt{\langle\phi|\Pi|\phi\rangle}}\nu\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}$
=$\stackrel{{\scriptstyle(a)}}{{=}}\max_{|\phi\rangle:|\phi\rangle=\Pi|\phi\rangle}\langle\phi|\nu|\phi\rangle\stackrel{{\scriptstyle(b)}}{{\leq}}\max_{|\phi\rangle}\langle\phi|\nu|\phi\rangle\stackrel{{\scriptstyle(c)}}{{=}}\|\nu\|_{\infty}.$
Here, $(a)$ follows from the fact that $\left\{\frac{\Pi|\phi\rangle}{\sqrt{\langle\phi|\Pi|\phi\rangle}}:|\phi\rangle\right\}=\left\{\phi:\Pi|\phi\rangle=|\phi\rangle\right\}$. $(b)$ is obtained because the restriction $\{|\phi\rangle:|\phi\rangle=\Pi|\phi\rangle\}$ is removed and $(c)$ is obtained from the definition of max-norm. ∎
Lemma D.2.
For any $\sigma\in\mathcal{S}(\mathcal{H})$ and projector $\mathrm{P}$ such that $\mathrm{P}\Pi_{\sigma}=\mathrm{P}$ then
$\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\frac{1}{\mathrm{Tr}\left(\mathrm{P}\sigma\right)}.$
with minimum is achieved at $\psi={\mathrm{P}}/{\mathrm{Tr}\left(\mathrm{P}\right)}$.
Proof.
We will prove this in 3 steps by showing the following as equal.
  1. 1.
    $\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}$
  2. 2.
    $\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}$
  3. 3.
    $\frac{1}{\mathrm{Tr}\left(\mathrm{P}\sigma\right)}$
1$\to$2: Observe that taking any $\tilde{c}>0$, we get
$\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\tilde{c}\psi}{\mathrm{Tr}\left(\tilde{c}\psi\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$
Now, note that $\mathrm{Tr}\left(\psi\sigma\right)$ is constant and so using Lemma D.1, we get $\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}$, so
$\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (71)
Substituting $\psi^{\prime}=\Pi_{\sigma}\psi\Pi_{\sigma}$, and from Theorem C.2 in Appendix C, we know that $\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}\subseteq\mathcal{P}(\mathrm{P})$. Hence,
$\min_{\psi^{\prime}\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}\leq\min_{\psi^{\prime}\in\{\Pi_{\sigma}\psi\Pi_{\sigma}:\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})\}}\left\|\frac{\psi^{\prime}}{\mathrm{Tr}\left(\psi^{\prime}\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\Pi_{\sigma}\psi\Pi_{\sigma}}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (72)
Hence
$\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (73)
But, from Theorem C.2 in Appendix C, we know that $\mathcal{P}(\mathrm{P})\subseteq\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, hence
$\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}\leq\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (74)
Combining the inequalities in (73) and (74), we get
$\min_{\psi\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}.$ (75)
2$\to$3: Taking $\phi=\frac{\psi}{\|\psi\|_{\infty}}$, we get the constrain as $\phi\in\mathcal{P}(\mathrm{P})$ and $\|\phi\|_{\infty}=1$ and minimum as
$\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\min_{\phi\in\mathcal{P}(\mathrm{P}),\|\phi\|_{\infty}=1}\frac{1}{\mathrm{Tr}\left(\phi\sigma\right)}=\left(\max_{\phi\in\mathcal{P}(\mathrm{P}),\|\phi\|_{\infty}=1}\mathrm{Tr}\left(\phi\sigma\right)\right)^{-1}.$
Focusing on $\phi$,
$\|\phi\|_{\infty}=1\Rightarrow\phi\leq\mathrm{I}\Rightarrow\mathrm{P}\phi\mathrm{P}\leq\mathrm{P}\Rightarrow\phi\leq\mathrm{P}\Rightarrow\sigma^{1/2}\phi\sigma^{1/2}\leq\sigma^{1/2}\mathrm{P}\sigma^{1/2}\Rightarrow\mathrm{Tr}\left(\phi\sigma\right)\leq\mathrm{Tr}\left(\mathrm{P}\sigma\right).$
Also, equality is obtained as $\phi=\mathrm{P}$ and so $\max_{\phi\in\mathcal{P}(\mathrm{P}),\|\phi\|_{\infty}=1}\mathrm{Tr}\left(\phi\sigma\right)=\mathrm{Tr}\left(\mathrm{P}\sigma\right)$. Thus we obtain,
$\min_{\psi\in\mathcal{P}(\mathrm{P})}\left\|\frac{\psi}{\mathrm{Tr}\left(\psi\sigma\right)}\right\|_{\infty}=\frac{1}{\mathrm{Tr}\left(\mathrm{P}\sigma\right)}.$
For $\psi\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P})$, we get equality at $\psi=\frac{\mathrm{P}}{\mathrm{Tr}\left(\mathrm{P}\right)}$. ∎
Lemma D.3.
For any $\rho,\sigma$ such that $\Pi_{\rho}=\Pi_{\sigma}$, $10<c_{r}<$ 1,
$\min_{\begin{subarray}{c}\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\\ \psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})\end{subarray}}\left\|\frac{c_{r}\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+\frac{(1-c_{r})\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}=\left(\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}\left(\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2},c_{r}\right)\right)^{-1}.$
Here, for any $\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0$, we have
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)\stackrel{{\scriptstyle\Delta}}{{=}}\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}.$
More detailed studey of the function is given in given in Appendix E.
Proof.
Following arguments similar to the previous proof. We get
$\min_{\begin{subarray}{c}\psi_{\max}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\max}),\\ \psi_{\min}\in\mathcal{S}(\mathrm{I}-\Pi_{\sigma}+\mathrm{P}^{\min})\end{subarray}}\left\|\frac{c_{r}\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+\frac{(1-c_{r})\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}$
$=\min_{\begin{subarray}{c}\psi_{\max}\in\mathcal{S}(\mathrm{P}^{\max}),\\ \psi_{\min}\in\mathcal{S}(\mathrm{P}^{\min})\end{subarray}}\left\|\frac{c_{r}\psi_{\max}}{\mathrm{Tr}\left(\psi_{\max}\sigma\right)}+\frac{(1-c_{r})\psi_{\min}}{\mathrm{Tr}\left(\psi_{\min}\sigma\right)}\right\|_{\infty}$
$=\min_{{\phi_{\max}\in\mathcal{S}(\mathrm{T}^{\max}),\phi_{\min}\in\mathcal{S}(\mathrm{T}^{\min})}}\left\|{c_{r}\sigma^{-1/2}\phi_{\max}\sigma^{-1/2}}+{(1-c_{r})\sigma^{-1/2}\phi_{\min}\sigma^{-1/2}}\right\|_{\infty}.$
Now, following on similar lines, we get
$=\min\left\{C:{c_{r}\sigma^{-1/2}\phi_{\max}\sigma^{-1/2}}+(1-c_{r}){\sigma^{-1/2}\phi_{\min}\sigma^{-1/2}}\leq C\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\ \text{for some }{\phi_{\max}\in\mathcal{S}(\mathrm{T}^{\max}),\phi_{\min}\in\mathcal{S}(\mathrm{T}^{\min})}\right\}$
=$=\left\{\max C:C\left(c_{r}\phi_{\max}+(1-c_{r})\phi_{\min}\right)\leq\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2}\ \text{for some }{\phi_{\max}\in\mathcal{S}(\mathrm{T}^{\max}),\phi_{\min}\in\mathcal{S}(\mathrm{T}^{\min})}\right\}^{-1}$
$=\left(\Upsilon_{\mathrm{T}^{\max},\mathrm{T}^{\min}}\left(\sigma^{1/2}\Pi_{\mathrm{P}^{\max}+\mathrm{P}^{\min}}\sigma^{1/2},c_{r}\right)\right)^{-1},$
thus completing the proof. ∎
Appendix E Definition and properties of $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)$
Definition 4.
Given a pair of orthogonal projectors $\Pi_{1}$ and $\Pi_{2}$ such that $\mathrm{Tr}\left(\Pi_{1}\Pi_{2}\right)=0$, we define a function $\Upsilon_{\Pi_{1},\Pi_{2}}:\mathcal{P}(\mathcal{H})\times\mathbb{R}\to\mathbb{R}$ as
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)=\{\max c:c(r\psi_{1}+(1-r)\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}.$
It denotes the largest possible trace of any matrix $\psi\in\mathcal{P}(\Pi_{1}+\Pi_{2})$ with property $\psi\leq\sigma$, $\mathrm{Tr}\left(\psi\Pi_{1}\right)=r\mathrm{Tr}\left(\psi\right)$ and $\Pi_{1}\psi\Pi_{2}=0$.
We will take cases when $\Pi_{1}+\Pi_{2}\leq\Pi_{\sigma}$. If this not the cases, $\Pi_{1}$ (and $\Pi_{2}$) can just be substituted by the projector corresponding to the subspace spanned by the intersection subspace of $\Pi_{1}$ (and $\Pi_{2}$) and $\Pi_{\sigma}$.
Properties:
  • $\Upsilon_{\Pi_{1},\Pi_{2}}(k\sigma,r)=k\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)$
  • $\{\max c:c(k_{1}\psi_{1}+k_{2}\psi_{2})\leq\sigma\text{ for some }\psi_{1}\in\mathcal{S}(\Pi_{1}),\psi_{2}\in\mathcal{S}(\Pi_{2})\}=\frac{1}{k_{1}+k_{2}}\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,\frac{k_{1}}{k_{1}+k_{2}}\right)$.
Some closed-form expressions: These closed form expression obtained for specific cases help in obtaining closed form expression in $\mathcal{C}3$ in Theorem III.4.
  • $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,0)=\mathrm{Tr}\left(\Pi_{\sigma^{-1/2}\Pi_{2}\sigma^{-1/2}}\sigma\right)=\mathrm{Tr}((\sigma^{-1/2}\Pi_{2}\sigma^{-1/2})^{-1})$
  • $\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,1)=\mathrm{Tr}\left(\Pi_{\sigma^{-1/2}\Pi_{1}\sigma^{-1/2}}\sigma\right)=\mathrm{Tr}((\sigma^{-1/2}\Pi_{1}\sigma^{-1/2})^{-1})$
  • If $\Pi_{\sigma}>\Pi_{1}+\Pi_{2}$, the substitution $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2},r\right)$ simplifies it as $\sigma^{1/2}\Pi_{\sigma^{-1/2}(\Pi_{1}+\Pi_{2})\sigma^{-1/2}}\sigma^{1/2}\in\mathcal{P}(\Pi_{1}+\Pi_{2})$. From here on-words, we can proceed with the assumption that $\Pi_{1}+\Pi_{2}=\Pi_{\sigma}$.
Lemma E.1.
If $\mathrm{Tr}\left(\Pi_{1}\right)=\mathrm{Tr}\left(\Pi_{2}\right)=1$, then
$\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=R_{\min}(\sigma,r\Pi_{1}+(1-r)\Pi_{2}).$
Proof.
If $\mathrm{Tr}\left(\Pi_{1}\right)=1,\psi_{1}\in\mathcal{S}(\Pi_{1})\Rightarrow\psi_{1}=\Pi_{1}$ and similarly, $\psi_{2}=\Pi_{2}$ and so,
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)=\{\max c:c(r\Pi_{1}+(1-r)\Pi_{2})\leq\sigma\}=R_{\min}(\sigma,r\Pi_{1}+(1-r)\Pi_{2}).$
Last step is obtained from the definition of $R_{\min}(\cdot,\cdot)$. ∎
Lemma E.2.
If $\sigma=\Pi_{1}\sigma\Pi_{1}+\Pi_{2}\sigma\Pi_{2}$ or say $\Pi_{1}\sigma\Pi_{2}=0$, then
$\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\begin{cases}\displaystyle\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r},&\displaystyle r\leq\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{\mathrm{Tr}\left(\sigma\right)},\vspace{1mm}\\ \displaystyle\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},&\text{else}.\end{cases}.$
Proof.
$\Upsilon_{\Pi_{1},\Pi_{2}}(\sigma,r)=\{\max c:c(r\Pi_{1}+(1-r)\Pi_{2})\leq\sigma\}.$
$c(r\psi_{1}+(1-r)\psi_{2})$ $\leq\sigma=\Pi_{1}\sigma\Pi_{1}+\Pi_{2}\sigma\Pi_{2}$
$\Rightarrow\begin{cases}cr\psi_{1}\leq\Pi_{1}\sigma\Pi_{1}&\Rightarrow cr\leq\mathrm{Tr}\left(\Pi_{1}\sigma\right)\\ \text{and}\\ c(1-r)\psi_{2}\leq\Pi_{2}\sigma\Pi_{2}&\Rightarrow c(1-r)\leq\mathrm{Tr}\left(\Pi_{2}\sigma\right)\end{cases}$
$\Rightarrow c\leq\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}.$
So, $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)\leq\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}$. Now, note that taking $\psi_{1}=\frac{\Pi_{1}\sigma\Pi_{1}}{\mathrm{Tr}\left(\Pi_{1}\sigma\right)},\psi_{2}=\frac{\Pi_{2}\sigma\Pi_{2}}{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}$, and $c=\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}$, we get
$c(r\psi_{1}+(1-r)\psi_{2})\leq\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}\left(r\frac{\Pi_{1}\sigma\Pi_{1}}{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}+(1-r)\frac{\Pi_{2}\sigma\Pi_{2}}{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}\right)\leq\Pi_{1}\sigma\Pi_{1}+\Pi_{2}\sigma\Pi_{2}=\sigma.$
Hence $\Upsilon_{\Pi_{1},\Pi_{2}}\left(\sigma,r\right)=\min\left\{\frac{\mathrm{Tr}\left(\Pi_{1}\sigma\right)}{r},\frac{\mathrm{Tr}\left(\Pi_{2}\sigma\right)}{1-r}\right\}$. ∎
Appendix F Proof of Theorem IV.1
We will use frequently here that $\mathrm{Tr}\left(\Gamma\nu\right)=0\Leftrightarrow\Gamma\in\mathcal{P}(\mathrm{I}-\Pi_{\nu})$ (See Lemma C.2 in Appendix C for proof). Starting with the definition of set $\mathcal{M}$ and the condition $e(\Lambda)=0$, the goal to find set of all ${\Lambda}$ such that $\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}$ and $e(\Lambda)=0$. So,
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\{\Lambda:e(\Lambda)=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}$
=$=\left\{\Lambda:\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}=0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (76)
The previous step follows from using the definition of $e(\Lambda)$. Now, for the fraction to be 00, numerator must be zero and denominator must remain non-zero. Hence, we have
$\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}=0$ $\Leftrightarrow\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0,\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))\neq 0$
$\Leftrightarrow\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0,p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0.$
Previous step is obtained by substituting $\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0$, and thus $\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))=p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$. Now using $\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)=0\Leftrightarrow\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0$, we get
$\frac{\mathrm{Tr}(p_{\sigma}\Lambda_{\rho}\sigma+p_{\rho}\Lambda_{\sigma}\rho)}{\mathrm{Tr}((\Lambda_{\rho}+\Lambda_{\sigma})(p_{\rho}\rho+p_{\sigma}\sigma))}=0$ $\Leftrightarrow\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0.$
Substituting it in (76), we get $\mathcal{E}_{s}(\rho,\sigma,p)=$
$\left\{\Lambda:\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}\geq 0,\Lambda_{\sigma}\geq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\right\}.$ (77)
Substituting $\mathrm{Tr}\left(\Lambda_{\rho}\sigma\right)=0,\Lambda_{\rho}\geq 0\Leftrightarrow\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})$ and similarly $\mathrm{Tr}\left(\Lambda_{\sigma}\rho\right)=0,\Lambda_{\sigma}\geq 0\Leftrightarrow\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho})$, we get
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\{\Lambda:\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\}.$
Now at max one of $p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)$ and $p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)$ can be 0 to ensure that $p_{\rho}\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)+p_{\sigma}\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0$, thus giving rise to one of the three cases, any POVM must satisfy as given below
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\left\{\Lambda:\begin{cases}\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\\ \text{ OR}\\ \Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\\ \text{ OR}\\ \Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\end{cases}\right\}.$
Note that $\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)=0\Leftrightarrow\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma})\Leftrightarrow\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ for the first case. Similarly $\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)=0\Leftrightarrow\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma})$ for the second case. Substituting these, we obtain
$\mathcal{E}_{s}(\rho,\sigma,p)$ =$=\left\{\Lambda:\begin{cases}\Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\qquad\text{ OR}\\ \Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho+\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}\qquad\text{ OR}\\ \Lambda_{\rho}\in\mathcal{P}(\mathrm{I}-\Pi_{\sigma}),\Lambda_{\sigma}\in\mathcal{P}(\mathrm{I}-\Pi_{\rho}),\mathrm{Tr}\left(\Lambda_{\rho}\rho\right)\neq 0,\mathrm{Tr}\left(\Lambda_{\sigma}\sigma\right)\neq 0,\Lambda_{\rho}+\Lambda_{\sigma}\leq\mathrm{I}.\end{cases}\right\}.$
Note that the set of measurements satisfying the first, second and third condition are $\mathcal{E}_{s}^{1}(\rho,\sigma)$, $\mathcal{E}_{s}^{2}(\rho,\sigma)$, $\mathcal{E}_{s}^{3}(\rho,\sigma)$ respectively. An arbitrary measurement should satisfy one of the three conditions, so the set of all such measurement is given by union of these three sets.